### v1

Abstract
Deep Generative Models (DGMs) have rapidly advanced in recent years, becoming essential tools in various fields due to their ability to learn complex data distributions and generate synthetic data. Their importance in transportation research is increasingly recognized, particularly for applications like traffic data generation, prediction, and feature extraction. This paper offers a comprehensive introduction and tutorial on DGMs, with a focus on their applications in transportation. It begins with an overview of generative models, followed by detailed explanations of fundamental models, a systematic review of the literature, and practical tutorial code to aid implementation. The paper also discusses current challenges and opportunities, highlighting how these models can be effectively utilized and further developed in transportation research. This paper serves as a valuable reference, guiding researchers and practitioners from foundational knowledge to advanced applications of DGMs in transportation research.
keywords:
[…]
Introduction
The rapid growth of artificial intelligence in transportation research over recent years has provided innovative solutions to address both longstanding and emerging transportation issues. Among the spectrum of artificial intelligence techniques, Deep Generative Models (DGMs) are increasingly gaining attention in transportation research due to their ability to learn the underlying patterns in large datasets and model the data distribution. Using the learned distributional characteristics, DGMs can generate synthetic data that accurately replicate real-world scenarios, estimate and predict agent-level trajectories, link-level and network-level traffic states, and provide latent representations of complex transportation systems. The primary purpose of this paper is to review the state-of-the-art theories and investigate the applications, potential benefits, and challenges of using DGMs in transportation research. Additionally, we aim to offer a comprehensive tutorial for researchers in transportation engineering who are interested in incorporating DGMs into their works. By providing a detailed overview of the current state-of-the-art models and literature with practical tutorials, this paper seeks to facilitate the adoption of DGMs and inspire further innovations in transportation research.
1.1 Background
Conventional deep learning models in transportation research have predominantly utilized discriminative modeling approaches. Discriminative models, as the name suggests, are designed to discriminate or differentiate patterns in data. These models learn a direct mapping from input variables ($\mathbf{x}$) to output variables ($\mathbf{y}$), effectively estimating the conditional probability $p(\mathbf{y}|\mathbf{x})$. In practice, this translates into the model’s ability to classify input data into predefined categories or predict target variables based on input features. For example, Convolutional Neural Networks (CNNs) and their variants are widely used to analyze spatial traffic patterns for classifying different traffic conditions; Recurrent Neural Networks (RNNs), including Long Short-Term Memory (LSTM) networks, model temporal dependencies for forecasting traffic variables over time; and Graph Neural Networks (GNNs) capture complex spatial relationships within transportation networks represented as graphs, facilitating tasks such as traffic prediction and network anomaly detection.
Even though discriminative models have achieved great success in transportation research, inherent limitations constrain their effectiveness. They require a structured form of extensive labeled (paired $\mathbf{x}$ and $\mathbf{y}$ data) datasets and are confined to the patterns present within the training data. Additionally, these models are unable to capture the underlying data distribution $p(\mathbf{x})$, which restricts their ability to understand complex patterns in the dataset fully. Consequently, these limitations hinder their capacity to generalize beyond the observed scenarios in the training dataset, i.e., the predictive ability is limited within what was given in the training dataset. Moreover, transportation data is inherently dynamic and complex, often characterized by uncertainty and rare events. Relying solely on discriminative models can thus impede the development of robust solutions capable of adapting to such complexities and predicting or generating rare events with non-zero probability.
Deep Generative Models (DGMs), on the other hand, are a class of machine learning models with the core objective of learning the joint probability $p(\mathbf{x},\mathbf{y})$ or the underlying data distribution $p(\mathbf{x})$. This capability enables DGMs to handle both discriminative tasks (such as classification and prediction) and generative tasks (such as data generation) within a unified framework. In contrast to traditional discriminative models, which are often limited by their reliance on extensive labeled datasets, DGMs can learn from the inherent patterns in data to generate realistic synthetic examples. In transportation research, DGMs can be used to create realistic traffic scenarios, such as time-space speed contour diagrams, simulate traveler behaviors under different conditions while accounting for the characteristics of individual agents, and generate synthetic datasets to augment rare events. Their ability to capture the complexities of traffic dynamics and traffic agent behavior directly addresses data scarcity issues and enhances the robustness and generalization of predictive models. This versatility makes DGMs an essential tool for managing the dynamic and uncertain nature of transportation systems.
From our viewpoint, the development of effective DGMs represents one of the most promising opportunities in modern transportation research. DGMs have the potential to revolutionize how we simulate, analyze, and optimize transportation systems by learning complex patterns from vast amounts of data without relying on strict assumptions or extensive model-based parameterization. While traditional theory-based models have been invaluable in understanding transportation dynamics, they often face challenges such as oversimplified assumptions and a narrow focus on specific scenarios. Despite these limitations, theory-based models remain essential, as they provide a foundation upon which data-driven methods like DGMs can build. Specifically, instead of replacing traditional approaches, DGMs serve as complementary and supplementary tools, enhancing and expanding upon established models. Their flexibility allows for the modeling of intricate dynamics and the generation of realistic, unseen scenarios, ultimately improving traffic simulations, scenario planning, and the implementation of digital twins in the transportation area.
Table 1: Comparison of previous survey papers
Reference Contribution Wide range of DGMs Transportation Application Tutorial Code
Harshvardhan et al. (2020) provide a comprehensive survey and implementation guide for machine and deep learning-based DGMs.
Bond-Taylor et al. (2021) provide a comparative review of DGMs, analyzing their strengths, weaknesses, and applications
De et al. (2022) provide a comprehensive review of DGMs in Industrial Internet of Things (IIoT) applications
Guo and Zhao (2022) provide systematic surveys of DGMS for graph generation with their applications, taxonomy, and analysis.
Lopez et al. (2020) provide a comprehensive review in the application of DGMs in molecular biology, highlighting their role in advancing scientific discoveries
Anstine and Isayev (2023) provide a comprehensive review in the application of chemical sciences with discussing their application challenges
Boquet et al. (2020) review the Variational Autoencoders (VAE) models to address key challenges in transportation
Lin et al. (2023a) provide comprehensive applications of Generative Adversarial Networks (GANs) in Intelligent Transportation Systems (ITSs)
Yan and Li (2023) provide a comprehensive reviews of DGMs in the ITS applications
Current Work provide a comprehensive review and implementation tutorial on DGMs focusing on their applications in transportation research
1.2 Relationship with Existing Surveys
DGMs have played a pivotal role in advancing machine learning over the past decade by enabling the generation and approximation of joint distributions for targets and training data. Several notable surveys have highlighted the significance and versatility of DGMs across various fields. For example, Harshvardhan et al. (2020) provided a comprehensive guide tracing the evolution from traditional machine learning-based generative models to advanced DGMs. Similarly, Bond-Taylor et al. (2021) reviewed a wide range of DGMs, including energy-based models, Variational Autoencoders (VAEs), autoregressive models, and normalizing flows, discussing their strengths, weaknesses, and applications. Area-specific surveys further illustrate the adaptability and impact of DGMs. For instance, De et al. (2022) explored DGMs’ applications in the Industrial Internet of Things (IIoT), while Guo and Zhao (2022) systematically reviewed DGMs for graph generation, focusing on their applications, taxonomy, and analysis. Additionally, Lopez et al. (2020) examined the role of DGMs in molecular biology, highlighting their utility in advancing scientific discoveries. Similarly, Anstine and Isayev (2023) provided a comprehensive review in the application of chemical sciences with discussing application challenges.
In transportation research, several review papers have similarly explored the potential of DGMs and highlighted their extensive applications in the field. For instance, Boquet et al. (2020) discussed the use of VAEs to address key transportation challenges, including traffic imputation, dimensionality reduction, and anomaly detection Similarly, Lin et al. (2023a) categorized the use of Generative Adversarial Networks (GANs) in autonomous driving, traffic flow research, and anomaly detection, while also identifying challenges and future research directions for integrating GANs into transportation research. However, these survey primarily focuses on typical models, with less coverage of other DGMs. Yan and Li (2023) systematically investigated the role of DGMs, or generative AI, in addressing key issues such as traffic perception, prediction, simulation, and decision-making within Intelligent Transportation Systems (ITS), while discussing existing challenges and future research directions. Even though this survey provides valuable insights, it would benefit from a more comprehensive explanation of fundamental DGMs and a deeper analysis of challenges and opportunities within the transportation context. Such enhancements could make DGMs more accessible and easier to adopt for a broader range of researchers in the field.
To address these gaps, this paper offers a comprehensive introduction and tutorial on DGMs in transportation applications. First, we provide a clear explanation of fundamental DGMs to ensure readers gain a solid understanding of core concepts. This is followed by a systematic review of state-of-the-art DGMs, with particular attention to their application in transportation research. We also include practical tutorial codes to guide researchers and practitioners in effectively applying these models to real-world transportation problems. Lastly, the paper concludes by emphasizing the importance of DGMs, discussing current research challenges, and exploring potential solutions, making it a valuable resource for those interested in exploring the use of DGMs in transportation.
1.3 Objectives of the Paper
This paper aims to provide a comprehensive survey of the application and potential of DGMs in the field of transportation research. The objectives of this survey paper are as follows:
  • To provide an accessible and comprehensive introduction to DGMs for transportation researchers. We aim to establish this paper as a starting point for those interested in exploring the potential of DGMs in transportation research.
  • To review the state-of-the-art in transportation research using DGMs, offering insights into the current landscape and practices across various transportation domains.
  • To contribute practical value by offering a tutorial section, providing hands-on guidance and resources for implementing DGMs in transportation research.
  • To identify and discuss the challenges, limitations, and opportunities of employing DGMs in transportation research.
With these objectives, we hope to contribute to the wider acceptance and understanding of DGMs within the field of transportation research. It is our belief that the adoption of these models can revolutionize the way we approach and solve transportation problems, leading to more robust, efficient, and sustainable systems.
1.4 Structure of Paper
In Section 2, we introduce the capabilities and applications of several key DGMs in the literature. Specifically, we introduce Variational Autoencoders (VAEs) in Section 2.3, Generative Adversarial Networks (GANs) in Section 2.4, Normalizing Flows (or flow-based generative models) in Section 2.5, Score-based generative models in Section 2.6 and Diffusion models in Section 2.7. In Section 3, we introduce state-of-the-art transportation research using DGMs. Particularly, we introduce applications of DGM 1) for generating realistic new data samples that can be applied in data synthesis, trajectory generation, and missing data imputation in Section 3.1, 2) for estimating and predicting distributions at three different levels of analyses in transportation research (agent-level, link-level, and region-level) in Section 3.2, and 3) for understanding underlying dynamics and learning unsupervised representations of data for applications like anomaly detection and mode choice analysis in Section 3.3. Readers primarily interested in the current practices can begin with Section 3 for a review of the latest literature in transportation research using DGMs, and then refer back to Section 2 for an introduction to the models themselves. In Section 4, we provide a tutorial that offers practical guidance on implementing DGMs in transportation research. We introduce two examples: 1) generating travel survey data in Section 4.1, and 2) generating highway traffic speed contour data in Section 4.2. In Section 5, we identify and discuss the challenges and opportunities associated with using DGMs in the transportation domain, emphasizing the importance of addressing these challenges for the successful adoption of DGMs. Finally, in Section 6, we summarize and conclude the paper.
Introduction to Deep Generative Models
2.1 Overview of Deep Generative Models
2.1.1   Background
Generative models are a class of machine learning models that aim to understand and capture the underlying probability distribution of a dataset. By learning this distribution, these models can generate new data points that are similar to those in the original dataset. Unlike discriminative models, which focus on modeling the conditional probability of the output given the input, $p(\mathbf{y}|\mathbf{x})$, generative models aim to learn the joint distribution $p(\mathbf{x},\mathbf{y})$ or $p(\mathbf{x})$. By understanding the joint distribution, generative models can also derive the conditional probability $p(\mathbf{y}|\mathbf{x})$ as $p(\mathbf{y}|\mathbf{x})=\frac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})}$. This means generative models can perform discriminative tasks effectively while also being capable of generating new data that aligns with the learned distribution. Therefore, generative models provide a broader framework that not only includes the functionalities of discriminative models but also extends beyond them to include data generation capabilities and model data generation process. Examples of earlier generative models include Gaussian Mixture Models, Hidden Markov Model, and Naive Bayes Classifier. However, they often rely on simplifying assumptions and may struggle with high-dimensional or complex data patterns.
Deep Generative Models (DGMs) build upon these concepts by leveraging deep learning architectures to model more complex, high-dimensional data without relying on such strong assumptions. By utilizing neural networks, DGMs can learn these complex, nonlinear relationships within data, allowing for more accurate modeling and generation of realistic data samples.
A key advantage of DGMs is their ability to generate data, which is essential in applications such as data augmentation, synthetic data creation, and data privacy enhancement. This capability allows the creation of new, synthetic datasets that can enhance the robustness and performance of machine-learning models, especially when real-world data is limited or sensitive. Another significant benefit of using DGMs is their capacity for probabilistic inference. It involves using the probability distributions learned by DGMs to do predictions and estimate uncertainties. The probabilistic inference is vital for understanding and modeling the uncertainties inherent in transportation systems, as it provides a range of possible outcomes along with their associated probabilities. Furthermore, DGMs are notable for their ability to extract deep insights from data and enable a wide range of applications through the manipulation and analysis of latent vectors. The latent vectors represent compressed, lower-dimensional forms of the input data that retain its essential features. This capability makes DGMs powerful tools for capturing and analyzing complex data relationships in transportation applications.
2.1.2   Preliminaries
Before exploring the specific types of generative models, it is important to grasp some fundamental concepts and terminology:
  • Bayes’ Rule: This is a fundamental theorem in probability theory that describes how to update the probability of a hypothesis based on new evidence. It combines prior knowledge (prior probability) with new evidence (likelihood) to form a posterior probability, which is a revised probability given the new information. Usually, in the context of DGMs, Bayes’ Rule is frequently used to derive an approximated loss function. The equation for Bayes’ Rule is:
    $p(A|B)=\frac{p(B|A)p(A)}{p(B)},$ (2.1)
    where $p(A|B)$ is the posterior probability of the hypothesis $A$ given the evidence $B$, $p(B|A)$ is the likelihood of observing the evidence $B$ given that $A$ is true, $p(A)$ is the prior probability of the hypothesis $A$ before observing the evidence, and $p(B)$ is the probability of observing the evidence $B$, also known as the marginal likelihood.
  • Data Distribution: This refers to the way data is spread or distributed in a particular space. Generative models aim to learn this distribution so they can generate new samples that resemble the original data.
  • Likelihood: Likelihood measures the plausibility of a model parameter value given specific observed data. In the context of generative models, it refers to how likely the observed data is under the model’s assumed data distribution.
  • KL divergence (Kullback–Leibler divergence): This is a measure of how one probability distribution is different from a second, reference probability distribution. It provides a way to quantify the difference between two probability distributions in bits. Mathematically, given two probability distributions $p(x)$ and $q(x)$ over the same random variable $x$, the KL divergence of $q$ from $p$ is defined as:
    $D_{KL}\left(p(\mathbf{x})||q(\mathbf{x})\right)=\mathbbm{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\log\left(\frac{p(\mathbf{x})}{q(\mathbf{x})}\right)\right]=\int_{\mathbf{x}\in\mathcal{X}}p(\mathbf{x})\log\left(\frac{p(\mathbf{x})}{q(\mathbf{x})}\right)dx,$ (2.2)
    where $p(\mathbf{x})$ and $q(\mathbf{x})$ are probability density functions (PDFs) over a continuous variable $\mathbf{x}$. The integral is taken over the entire range of $\mathbf{x}$ where the distributions are defined. It is worth noting that KL divergence is non-negative and is not symmetric, meaning $D_{KL}(p(\mathbf{x})||q(\mathbf{x}))\neq D_{KL}(q(\mathbf{x})||p(\mathbf{x}))$.
  • Latent Space: This is a conceptual space where the high-dimensional data is compressed into lower dimensions. The data in this latent space is usually the input to the generative part of the model that produces the output samples.
2.1.3   Model Training Objective
The objective of training DGMs is to minimize the discrepancy between the true data distribution $p_{\text{data}}(\mathbf{x})$ and the model’s estimated distribution $p_{\text{model}}(\mathbf{x;\theta})$. This discrepancy can be expressed using the KL divergence. The goal is to adjust the model parameters $\theta$ such that the model distribution closely approximates the true data distribution. This is expressed in the optimization problem:
$\begin{split}\theta^{*}=\arg\min_{\theta}D_{KL}(p_{\text{data}}(\mathbf{x})||p_{\text{model}}(\mathbf{x};\theta)).\end{split}$ (2.3)
However, in practice, we do not have direct access to $p_{\text{data}}(\mathbf{x})$, and hence we rely on an empirical dataset (or training dataset) to represent the true data distribution. Given that direct computation of KL divergence is not feasible without knowledge of $p_{\text{data}}(\mathbf{x})$, DGMs typically rely on maximizing the likelihood of the observed data under the model. This approach aligns with the principle of maximum likelihood estimation (MLE), which posits that the optimal model parameters $\theta$ are those that maximize the likelihood of the training data appearing under the model’s assumed probability distribution. Mathematically, this is represented as:
$\begin{split}\theta^{*}=\arg\max_{\theta}\prod_{i=1}^{N}p_{\text{model}}(\mathbf{x}_{i};\theta),\end{split}$ (2.4)
where $\mathbf{x}_{i}$ denotes an instance of the training data. To simplify calculations and enhance numerical stability, we typically work with the logarithm of the likelihood function, transforming the product into a summation. This is referred to as the log-likelihood, and it modifies our optimization problem to:
$\begin{split}\theta^{*}=\arg\max_{\theta}\sum_{i=1}^{N}\log p_{\text{model}}(\mathbf{x}_{i};\theta),\end{split}$ (2.5)
which is essentially the mean log-likelihood over all $N$ data points in the training dataset. This is the empirical estimate of the expected log-likelihood under the data distribution $p_{\text{data}}(x)$:
$\begin{split}\theta^{*}&=\arg\max_{\theta}\mathbbm{E}\left[\log p_{\text{model}}(\mathbf{x};\theta)\right]\\ &=\arg\max_{\theta}\mathbbm{E}\left[\log p_{\theta}(\mathbf{x})\right].\end{split}$ (2.6)
2.2 DGM Classification
DGMs can be classified into different categories based on their approach to maximizing data likelihood. The first one is explicit density models that directly model the probability distribution $p_{\theta}(\mathbf{x})$. These models use the likelihood function during training to align the generated data distribution with the true data distribution. In contrast, implicit density models do not explicitly use $p_{\theta}(\mathbf{x})$ during training. Instead, they employ alternative methods that theoretically achieve the goal of maximizing data likelihood by implicitly capturing the data distribution. Explicit density models can be further categorized into tractable density models and intractable density models based on whether the likelihood computation during training is directly calculable or must be approximated.
2.2.1   Explicit Tractable Density Models
Explicit Tractable Density Models perform direct and explicit computations of the likelihood, making them highly interpretable and theoretically robust. Examples include:
  • Autoregressive Models: Examples include PixelRNN (Van Den Oord et al., 2016), PixelCNN (Van den Oord et al., 2016), and NADE (Uria et al., 2016). In these models, each output is generated sequentially with each output conditioned on previous outputs. This sequential generation ensures that the likelihood of each output can be directly calculated.
  • Normalizing Flows: These models utilize invertible transformations to map complex data distributions to simpler, known parametric distributions (such as Gaussian distribution). The invertibility of these transformations allows for the exact computation of the likelihood of any given input, making these models particularly powerful for tasks requiring detailed density estimation.
2.2.2   Explicit Intractable Density Models
Explicit Intractable Density Models aim to directly model $p_{\theta}(\mathbf{x})$, but they rely on approximation methods due to the infeasibility of exact computation in complex data scenarios. These models include:
  • Variational Autoencoders (VAEs): VAEs use an encoder-decoder architecture where the encoder approximates the posterior distribution of latent variables, and the decoder approximates the data distribution conditioned on these latent variables. The likelihood is indirectly maximized through a lower bound estimate, known as the Evidence Lower Bound (ELBO).
  • Diffusion Models: These models gradually convert data into a known noise distribution and learn to reverse this process. Training involves approximating the reverse of this diffusion process, which indirectly maximizes the data likelihood.
  • Score-Based Generative Models: These models learn the score (gradient of log probability) of the data distribution and use it to generate samples by iteratively refining noise samples. The model indirectly maximizes the likelihood through the score-matching technique.
2.2.3   Implicit Density Models
Implicit Density Models do not explicitly use $p_{\theta}(\mathbf{x})$ during training. Instead, they employ alternative approaches that can theoretically align with the goal of maximizing data likelihood by implicitly capturing the data distribution.
  • Generative Adversarial Networks (GANs): GANs consist of two competing networks—a generator and a discriminator. The generator produces samples aimed at being indistinguishable from real data, while the discriminator evaluates their authenticity. Even though $p_{\theta}(\mathbf{x})$ is not explicitly computed, the adversarial training process theoretically leads to the generator capturing the true data distribution, resulting in implicitly maximizing the data likelihood.
2.3 Variational Autoencoder
(a) Model architecture of Autoencoder (AE)
(b) Model architecture of Variational Autoencoder (VAE)
Figure 1: Schematic overview of AE and VAE
Variational Autoencoders (VAEs) are a class of DGMs that combines the structure of Autoencoders with variational inference. Since the introduction in Kingma and Welling (2013), VAEs have significantly influenced the overall field of generative models. Autoencoders (AEs) typically comprise two main components: an encoder and a decoder. AEs were first introduced by Hinton and Salakhutdinov (2006) for data compression and dimensionality reduction. Specifically, AEs compress input data $\mathbf{x}$ into a lower-dimensional latent representation $\mathbf{z}$ using the $\phi$-parameterized encoder $f_{\phi}(\mathbf{x})=\mathbf{z}$ and reconstruct the input data from the latent representation with the $\theta$-parameterized decoder $g_{\theta}(\mathbf{z})=\mathbf{x}$. The objective function of an AE is typically:
$\begin{split}\mathcal{L}_{AE}=\|\mathbf{x}-\mathbf{x^{\prime}}\|^{2}=\|\mathbf{x}-g_{\theta}(f_{\phi}(\mathbf{x}))\|^{2}\end{split},$ (2.7)
where $\mathbf{x^{\prime}}$ is the reconstructed data of $\mathbf{x}$.
In contrast, the primary objective of VAEs is data generation, focusing on training a high-performing decoder. The decoder samples from the latent space to generate new data that resemble the original training dataset. Unlike traditional AEs that learn deterministic functions, $f_{\phi}$ and $g_{\theta}$, VAE model probabilistic (or generative) encoder and decoder, $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $p_{\theta}(\mathbf{x}|\mathbf{z})$, by introducing intermediate latent variable $z$ and using variational inference. The $\phi$-parameterized encoder, represented by $q_{\phi}(\mathbf{z}|\mathbf{x})$, transforms input data $\mathbf{x}$ into a latent representation $\mathbf{z}$ in a lower-dimensional latent space. Notably, instead of learning a deterministic mapping to a single latent vector, VAE’s encoder estimates the parameters of a probability distribution in the latent space. The $\theta$-parameterized decoder, represented by $p_{\theta}(\mathbf{x}|\mathbf{z})$, takes the encoded latent vector $\mathbf{z}$ and reconstructs the original data point $\mathbf{x}$.
The probability $p_{\theta}(\mathbf{x})$ with the latent vector $\mathbf{z}$ can be formulated as:
[…]
However, Equation (2.8) is intractable due to the integral over the $\mathbf{z}$ space. Even though Monte Carlo methods can be used to approximate the integral and apply maximum likelihood estimation, they may result in suboptimal generation due to poor-quality samples being assigned higher likelihoods. To address this issue, VAEs incorporate the encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ into their objective function, reformulated as:
$\begin{split}\log p(\mathbf{x})&=\mathbbm{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p(\mathbf{x})\right]\\ &=\mathbbm{E}_{\mathbf{z}}\left[\log\frac{p_{\theta}(\mathbf{x},\mathbf{z})}{p_{\theta}(\mathbf{z}|\mathbf{x})}\right]&&\Rightarrow\textit{ \small Bayes' Rule}\\ &=\mathbbm{E}_{\mathbf{z}}\left[\log\frac{p_{\theta}(\mathbf{x}|\mathbf{z})p(\mathbf{z})}{p_{\theta}(\mathbf{z}|\mathbf{x})}\right]\\ &=\mathbbm{E}_{\mathbf{z}}\left[\log\frac{p_{\theta}(\mathbf{x}|\mathbf{z})p(\mathbf{z})}{p_{\theta}(\mathbf{z}|\mathbf{x})}\frac{q_{\phi}(\mathbf{z}|\mathbf{x})}{q_{\phi}(\mathbf{z}|\mathbf{x})}\right]&&\Rightarrow\textit{ \small Multiply by $\frac{q_{\phi}(\mathbf{z}|\mathbf{x})}{q_{\phi}(\mathbf{z}|\mathbf{x})}=1$}\\ &=\mathbbm{E}_{\mathbf{z}}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-\mathbbm{E}_{\mathbf{z}}\left[\log\frac{q_{\phi}(\mathbf{z}|\mathbf{x})}{p(\mathbf{z})}\right]+\mathbbm{E}_{\mathbf{z}}\left[\log\frac{q_{\phi}(\mathbf{z}|\mathbf{x})}{p_{\theta}(\mathbf{z}|\mathbf{x})}\right]&&\Rightarrow\textit{ \small Logarithms}\\ &=\mathbbm{E}_{\mathbf{z}}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)+D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p_{\theta}(\mathbf{z}|\mathbf{x})\right).\end{split}$ (2.9)
The first term in Equation (2.9), $\mathbbm{E}_{z}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]$, is the likelihood of $\mathbf{x}$ generated from the decoder based on the sampled $\mathbf{z}$ from the posterior distribution. This term can be estimated through sampling using the reparameterization trick, which will be discussed later. The second term is the KL divergence between the approximate posterior distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ and the prior distribution $p(\mathbf{z})$. Assuming both distributions follow a tractable distribution (typically Gaussian), this term can be computed in closed form. The third term is intractable because the true posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x})$ involves evaluation of intractable Equation (2.9) $\left(p_{\theta}(\mathbf{z}|\mathbf{x})=\frac{p_{\theta}(\mathbf{x}|\mathbf{z})p(\mathbf{z})}{{p_{\theta}(\mathbf{x})}}\right)$. However, the third term is always non-negative due to the properties of KL divergence. Therefore, a new objective function called the Evidence Lower BOund (ELBO), $\mathcal{L}(\mathbf{x};\theta,\phi)$, can be defined as:
$\log p_{\theta}(\mathbf{x})\geq\mathbbm{E}_{\mathbf{z}}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)=\mathcal{L}(\mathbf{x};\theta,\phi).$ (2.10)
Equation (2.10) demonstrates that the ELBO is a ‘lower bound’ to the log-likelihood of the data. Therefore, maximizing the ELBO with respect to the parameters of the encoder and decoder, $\phi$ and $\theta$, also maximizes the true log-likelihood ($\log p_{\theta}(\mathbf{x})$). The ELBO can be efficiently estimated using stochastic gradient-based optimization methods without involving the intractable posterior $p_{\theta}(\mathbf{z}|\mathbf{x})$. Finally, the maximizing objective function of VAE can be expressed as:
$\mathcal{L}(\mathbf{x};\theta,\phi)=\mathbbm{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right),$ (2.11)
where the first term, $\mathbbm{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]$, is called reconstruction error, which measures the VAE’s capability to accurately reconstruct the input data from its latent variable generated by the encoder network. The second term, $D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)$, is called regularization, which focuses on aligning the distribution of the latent variables with the assumed distribution.
A key challenge in training VAEs is estimating the gradient of the expected value term in the ELBO to the encoder network parameters $\phi$. This involves differentiating through the random sampling operation involved in generating the latent variable $\mathbf{z}$, which is inherently non-deterministic and does not permit direct differentiation. The reparameterization trick addresses this challenge by reparameterizing the random variable $\mathbf{z}$ such that the randomness is independent of the parameters. This enables the model to backpropagate gradients through the deterministic part of the reparameterization, while the stochasticity is handled separately. Instead of sampling $\mathbf{z}$ directly from $q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mu(\mathbf{x};\phi),\sigma(\mathbf{x};\phi)^{2})$, the model samples a standard Gaussian noise variable $\epsilon\sim\mathcal{N}(0,I)$ and reparameterizes $\mathbf{z}$ as follows:
$\mathbf{z}=\mu(\mathbf{x};\phi)+\sigma(\mathbf{x};\phi)\odot\epsilon,$ (2.12)
where $\odot$ denotes the element-wise multiplication. The gradients with respect to $\phi$ can be computed since the sampling operation is deterministic for $\phi$, and all randomness is relegated to $\epsilon$.
Conditional VAEs (CVAEs) (Sohn et al., 2015) is one of the extensions of the standard VAEs. The core innovation of CVAEs lies in their ability to incorporate additional conditional information, often in the form of labels or related data, directly into the generative process. This conditioning allows the model to generate more targeted and context-specific data. Unlike traditional VAEs, CVAEs leverage the extra conditional information to produce outputs that are not only high in quality but also relevant to the specified conditions. This is achieved by modifying both the encoder and decoder to accept and process the conditional information alongside the input data. As a result, CVAEs have a degree of control and specificity in the generated output, such as generating images of a certain class or style, customizing text generation, and enhancing recommendation systems. They also offer benefits in interpretability and the potential for disentangled representation learning. The CVAE framework has thus emerged as a powerful tool for controlling the generated output and improving the model performance. Applications of CVAE in transportation research varies from traffic data generation, prediction and classification.
Adversarial Autoencoders (AAEs) (Makhzani et al., 2015) represent an innovative fusion of VAEs (VAEs) and Generative Adversarial Networks (GANs; See Section 2.4 for more details), designed to enhance the capabilities of generative models. The primary advantage of AAEs lies in their ability to impose arbitrary prior distributions on the latent space, going beyond the Gaussian priors typically used in VAEs. This flexibility allows for the modeling of complex data distributions more effectively. AAEs consist of a standard autoencoder architecture, complemented by an adversarial network that enforces the latent space to conform to the chosen prior distribution. To be more specific, the fundamental condition in VAE loss is that KL divergence between the distribution of encoded input data and prior can be calculated. The function of KL divergence in loss term is making the distribution of encoded input data ($q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)$) the same as the prior distribution $p(\mathbf{z})$, which is the same logic of GAN. Therefore, AAE utilizes the discriminator term of GAN to replace the KL divergence of VAE, making it possible to use arbitrary prior distribution. Through adversarial training, AAEs generate sharper, more detailed outputs compared to traditional VAEs, especially noticeable in tasks like image generation. They are versatile in applications, ranging from semi-supervised learning to unsupervised clustering and anomaly detection. Additionally, AAEs offer better control over the generation process, including conditional generation and the learning of disentangled representations, making them suitable for tasks requiring precise control and interpretability. The integration of adversarial principles into autoencoders has thus positioned AAEs as a powerful and flexible tool in the realm of generative modeling, addressing key challenges of traditional VAEs and opening new avenues in machine learning research.
2.4 Generative Adversarial Networks
Generative Adversarial Networks (GANs) are another class of generative models, introduced in Goodfellow et al. (2014). GANs have attracted considerable attention due to their capability to generate highly realistic data, and they have been widely used in a wide range of applications, including image synthesis, text generation, and data augmentation. Unlike other types of DGMs, GANs do not explicitly model data distribution or data likelihood. Instead, they employ a game-theoretical approach to train a model focused on generating realistic data.
Figure 2: Schematic overview of Generative Adversarial Network (GAN)
GANs consist of two primary components: a Generator (G) and a Discriminator (D). The Generator’s task is to produce realistic synthetic data, while the Discriminator focuses on distinguishing between real data and synthetic data generated by the Generator. This competitive dynamic resembles a two-player min-max game, where the Generator attempts to deceive the Discriminator, and the Discriminator endeavors to avoid being deceived. Formally, let the parameters of the Generator $G$ be denoted as $\theta$ and the parameters of the Discriminator $D$ be denoted as $\phi$. The Generator $G$ takes as input a random noise vector $\mathbf{z}$, sampled from a known distribution (often a standard normal distribution), and transforms it into a data instance $\mathbf{x}=G_{\theta}(\mathbf{z})$. The Generator is similar to the decoder of VAE in that both aim to learn the mapping from random noise to the data sample. The Discriminator $D$ takes as input a data instance, which can either be real data $\mathbf{x}$ or generated data $G_{\theta}(\mathbf{z})$, and outputs a scalar representing the probability that the input data is real, denoted as $D_{\phi}(\mathbf{x})$. The learning process is guided by a value function $V(\phi,\theta)$, defined as:
$V(\phi,\theta)=\mathbbm{E}_{\mathbf{x}\sim p_{\text{data}}(\mathbf{x})}[\log D_{\phi}(\mathbf{x})]+\mathbbm{E}_{\mathbf{z}\sim p(\mathbf{z})}[\log(1-D_{\phi}(G_{\theta}(\mathbf{z})))],$ (2.13)
where $p_{\text{data}}$ represents the true data distribution, and $p({\mathbf{z}})$ denotes the distribution of the input noise vectors. The first and second terms in the value function correspond to the log probability of the Discriminator correctly identifying real and generated data, respectively. The learning process involves finding optimal parameters, $\phi^{*}$ and $\theta^{*}$, for both $D$ and $G$, through gradient-based methods as shown below. This process alternates between the following two steps:
$\phi^{*},\theta^{*}=\min_{{\theta}}\max_{{\phi}}V(\phi,\theta),$ (2.14)
  1. 1.
    Maximizing $V(\phi,\theta)$ with respect to $\phi$: This step enhances the Discriminator’s capability to distinguish between real and generated data.
  2. 2.
    Minimizing $V(\phi,\theta)$ with respect to $\theta$: This step improves the Generator’s ability to generate synthetic data that can deceive the Discriminator.
This process does not explicitly involve the evaluation of the likelihood function; however, it implicitly maximizes the data likelihood and minimizes the difference between true data distribution $p_{\text{data}}$ and model distribution $p_{\text{model}}$. Consider $p_{\text{model}}$ as the learned model distribution imposed by the $\mathbf{z}\sim p(\mathbf{z})$ and the $\theta$-parameterized mapping function $G_{\theta}$. If $G_{\theta}$ is fixed, then $D$ would converge to $D_{\phi}^{*}(\mathbf{x})=\frac{p_{\text{model}}(\mathbf{x})}{p_{\text{data}}(\mathbf{x})+p_{\text{model}}(\mathbf{x})}$, since
[…]
which is optimal when $D_{\phi}^{*}(\mathbf{x})=\frac{p_{\text{model}}(\mathbf{x})}{p_{\text{data}}(\mathbf{x})+p_{\text{model}}(\mathbf{x})}$.
For the optimal discriminator, the objective of the generator can be expressed as
[…]
where $D_{JS}$ is the Jensen-Shannon divergence. The minimization is achieved when $D_{JS}\left(p_{\text{model}}(\mathbf{x})||p_{\text{data}}(\mathbf{x})\right)$ is minimized, i.e., the model distribution $p_{\text{model}}$ is close to the data distribution $p_{\text{data}}$.
The training process continues until a balance between the Generator and the Discriminator is achieved, where the Generator can produce realistic data and the Discriminator cannot differentiate between real and generated data, i.e., $D^{*}(\mathbf{x})=\frac{p_{\text{model}}(\mathbf{x})}{p_{\text{data}}(\mathbf{x})+p_{\text{model}}(\mathbf{x})}=\frac{p_{\text{data}}(\mathbf{x})}{p_{\text{data}}(\mathbf{x})+p_{\text{data}}(\mathbf{x})}=0.5$, for all $\mathbf{x}$.
However, in practice, achieving equilibrium is challenging due to issues like mode collapse, vanishing gradients, and non-convergence. Various modifications and extensions of the original GAN framework have been proposed to mitigate these issues, such as Wasserstein GANs (Arjovsky et al., 2017), Least-Squares GANs (Mao et al., 2017), and Conditional GANs (Shahbazi et al., 2022), each offering unique perspectives and solutions. Despite these challenges, the ability to generate highly realistic and diverse data has made GAN-based models essential tools in many domains of machine learning and data science.
Conditional GANs (Mirza and Osindero, 2014), commonly known as cGANs, are an advanced adaptation of the traditional GAN framework. The primary motivation behind cGANs is to give more control over the data generation process. While traditional GANs are adept at generating data from a learned distribution, cGANs add an additional layer of conditioning, usually based on labels or external information. For example, in the case that the user wants to generate two distinctive classes of outputs, standard GANs would require two separate models to learn each distribution or, at best, generate a mix without explicit control over the output. On the other hand, cGANs allow the model to generate images of a specific category based on a label provided to it. The beauty of cGANs lies in the fact that both the generator and the discriminator are conditioned on this label. The generator utilizes this label to produce data, while the discriminator uses it to discern the authenticity of the generated data. The mathematical representation encapsulates this conditioning through additional input layers for the generator and discriminator. When producing an output sample, the generator does not merely rely on the random noise vector but also factors in the label, ensuring the generated output aligns with the desired category. On the flip side, the discriminator evaluates both the output and its corresponding label, ensuring that fake (generated) data is not only realistic but is also consistent with its associated type.
Generative Adversarial Imitation Learning (Ho and Ermon, 2016), or GAIL, is an innovative fusion of imitation learning principles with the GAN framework. Imitation learning, at its core, revolves around the idea of learning a policy (or behavior) by using experts’ demonstrations. GAIL captures this essence but employs the adversarial training approach of GANs to achieve it. In the GAIL framework, the generator is no longer producing static data; instead, it produces sequences of actions aiming to imitate an expert’s behavior. Meanwhile, the discriminator’s role is to distinguish between sequences produced by the generator and those exhibited by the expert. Traditional imitation learning often relies on methods like behavioral cloning, where the agent directly learns from expert trajectories. However, this method is prone to compounding errors and struggles with situations not encountered during training. GAIL, on the other hand, captures the underlying reward structure by using the adversarial training framework.
2.5 Normalizing Flows (Flow-based Generative Models)
A flow-based deep generative model learns the likelihood by using normalizing flows. A normalizing flow describes a method for constructing a complex distribution by using a series of invertible mapping functions (Rezende and Mohamed, 2015). The idea behind using normalizing flows to learn the likelihood of the real data is that a complex distribution can be learned by sequentially applying multiple normalizing flows to a simple base distribution (e.g., Gaussian distribution).
For a multivariate vector x, the objective of normalizing flow is to learn the probability density function $p(\textbf{x})$. The density $p(\textbf{x})$ is transformed into a simple distribution $p(\mathbf{z}_{0})$ (independent multivariate Gaussian distribution). This transformation $f$ is composed of $K$ invertible (bijective) functions $f_{i},i\in\{1,...K\}$. This transformation can be denoted as follows:
$\begin{split}&\mathbf{x}=\mathbf{z}_{K}=f_{K}\circ f_{K-1}\circ\cdots\circ f_{1}(\mathbf{z}_{0})=f(\mathbf{z}_{0}),\\ \end{split}$ (2.17)
where we denote $\mathbf{z}_{i}$ as latent vector after applying $i$ invertible functions ($f_{1},\cdots,f_{i}$).
what is the meaning of $\circ$
Figure 3: Schematic overview of normalizing flow
By definition,
$\begin{split}&\mathbf{z}_{i}=f_{i}(\mathbf{z}_{i-1}),\text{ thus }\mathbf{z}_{i-1}=f_{i}^{-1}(\mathbf{z}_{i})\\ \end{split}.$ (2.18)
The log-likelihood of probability density function of $i$-th latent vector, $p(\mathbf{z}_{i})$, can be calculated as follows:
[…]
As a results, the probability density of $\mathbf{x}$ can be calculated as follows:
$\begin{split}\log p(\mathbf{x})=\log\pi_{K}(\mathbf{z}_{K})&=\log p_{K-1}(\mathbf{z}_{K-1})-\log\left|\det\left({\frac{df_{K}}{dz_{K-1}}}\right)\right|\\ &=\log p_{K-2}(\mathbf{z}_{K-2})-\log\left|\det\left({\frac{df_{K-1}}{dz_{K-2}}}\right)\right|-\log\left|\det\left({\frac{df_{K}}{dz_{K-1}}}\right)\right|\\ &=\cdots\\ &=\log p_{0}(z_{0})-\sum_{i=1}^{K}\log\left|\det\left({\frac{df_{i}}{dz_{i-1}}}\right)\right|.\end{split}$ (2.20)
Finally, the objective of training for a generative model is maximizing the likelihood of a training set, $\mathcal{D}$, as follows:
$\begin{split}\mathcal{L}(\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\log p(\mathbf{x})\end{split}.$ (2.21)
In practice, there are two conditions to consider when deciding proper functions for normalizing flows. The first condition is that by definition, the functions should be invertible. Also, computing the Jacobian determinant should be feasible since usually computing the Jacobian of functions and computing the determinant are both computationally expensive. As a result, properly defining the function $f$ is the key to normalizing flows.
There are several feasible functions from previous studies that satisfy the constraints: linear function (Rezende and Mohamed, 2015), $1\times 1$ convolution (Kingma and Dhariwal, 2018), and affine coupling layer (Dinh et al., 2014; Durkan et al., 2019). In this paper, we introduce the affine coupling layer as an example since this is one of the widely used forms due to its ability to deal with high-dimensional data such as image and sound. The affine coupling layer from Dinh et al. (2016) can be formulated as follows:
$\begin{split}&\mathbf{y}_{1:d}=\mathbf{x}_{1:d}\\ &\mathbf{y}_{d+1:D}=\mathbf{x}_{d+1:D}\odot\exp{\left(s(\mathbf{x}_{1:d})\right)}+t(\mathbf{x}_{1:d}),\\ \end{split}$ (2.22)
where $\mathbf{y}_{1:d}$ is the first $d$ elements of $\mathbf{y}$, $\mathbf{y}_{d+1:D}$ is the rest of $\mathbf{y}$, $\mathbf{x}_{1:d}$ is the first $d$ elements of $\mathbf{x}$, $\mathbf{x}_{d+1:D}$ is the rest of $\mathbf{x}$, $s$ and $t$ are scale and translation functions, which are usually implemented as multi-layer perceptrons (MLP), and $\odot$ is the element-wise multiplication operator. As noted in Dinh et al. (2016), the affine coupling layer is an invertible function and guarantees fast computation of the determinant of the Jacobian matrix.
First, the inverse function of Equation (2.22) is as follows:
$\begin{split}&\mathbf{x}_{1:d}=\mathbf{y}_{1:d}\\ &\mathbf{x}_{d+1:D}=\left(\mathbf{y}_{d+1:D}-t\left(\mathbf{y}_{1:d}\right)\right)\odot\exp{\left(-s\left(\mathbf{y}_{1:d}\right)\right)}.\\ \end{split}$ (2.23)
Also, the Jacobian matrix of this transformation is
$\begin{split}\frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}=\begin{bmatrix}\mathbbm{I}_{d}&0\\ \frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}&\operatorname{diag}\left(\exp{\left[s(\mathbf{x}_{1:d})\right]}\right)\end{bmatrix},\end{split}$ (2.24)
where $\mathbbm{I}_{d}$ is a $d\times d$ identity matrix, $\operatorname{diag}(*)$ is the diagonal matrix whose diagonal entries correspond to the given matrix and non-diagonal entries are zero.
As a result, the Jacobian matrix shown in Equation (2.24) is a lower-triangular matrix. Therefore, the determinant of the Jacobian matrix can be calculated as follows:
$\begin{split}\left|\det\left(\frac{\partial y}{\partial x^{\top}}\right)\right|&=\prod_{i=1}^{D}\left[\frac{\partial y}{\partial x^{\top}}\right]_{i,i}=\prod_{i=1}^{d}\exp{\left[s(\mathbf{x}_{1:d})\right]_{i,i}}\\ &=\exp{\left(\sum_{i=1}^{d}\left[s(\mathbf{x}_{1:d})\right]_{i,i}\right)},\\ \end{split}$ (2.25)
and
$\begin{split}\log\left|\det\left(\frac{\partial y}{\partial x^{\top}}\right)\right|&=\sum_{i=1}^{d}\left[s(\mathbf{x}_{1:d})\right]_{i,i},\\ \end{split}$ (2.26)
where $[*]_{i,i}$ is $i$-th diagonal entry of the given matrix.
2.6 Score-Based Generative Models
Score-based generative models are a class of generative models that leverage the concept of “score” of a probability density function to generate new data samples. The term score is used in a statistical context, referring to the gradient (or intuitively, slope) of the log-likelihood function with respect to the data, $\mathbf{x}$. Formally, the score is denoted as $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Intuitively, the score tells us how to change $\mathbf{x}$ to increase the probability of $\mathbf{x}$ under the model’s current estimate of the probability density function. A higher score indicates a direction in which the data point $\mathbf{x}$ could be adjusted to make it more likely under the model.
Figure 4: Schematic overview of score-based generative model
In score-based generative models, this concept is used to iteratively refine samples from an initial noise distribution, guiding them towards regions of higher probability under the target distribution (i.e., the distribution of the training data). This is achieved by estimating the score function at various points in the data space and using it to perform gradient ascent on the log-likelihood landscape. Essentially, by following the direction of the score, new data samples can be generated that are likely under the model’s learned distribution, thus resembling the characteristics of the training data.
The goal of (deep) generative models is to learn the underlying data distribution $p(\mathbf{x})$ given training data samples drawn from true data distribution, $\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{N}\}$ with $\mathbf{x}_{i}\sim p(\mathbf{x}),\forall i\in{1,\cdots,N}$. The score-based generative models (Song and Ermon, 2019; Song et al., 2020b) focuses on the development of a “score network,” denoted as $s_{\theta}(\mathbf{x})$, which is essentially a $\theta$-parameterized neural network designed to approximate the score of the data distribution; that is, $s_{\theta}(\mathbf{x})$ aims to be a close estimate of $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. The score network can be trained by minimizing the squared $L2$ distance between the ground-truth score and the estimated score by the score network as follows:
$\begin{split}\mathbbm{E}_{\mathbf{x}\sim p(\mathbf{x})}\|\nabla_{\mathbf{x}}\log p(\mathbf{x})-s_{\theta}(\mathbf{x})\|^{2}_{2}.\end{split}$ (2.27)
A significant challenge with this approach is that it involves direct access to $\nabla_{\mathbf{x}}\log p(\mathbf{x})$, which is not practically feasible. To address this issue, a substantial body of research has been dedicated to developing alternative methodologies, collectively known as “score matching” (Hyvärinen and Dayan, 2005; Vincent, 2011; Song et al., 2020a) that minimizes Equation (2.27) without having to have access to the ground-truth data score.
One of the main concerns in score-based generative models is figuring out how the score network can be efficiently and accurately trained. Song and Ermon (2019) identified several challenges in training the score network. One of the main challenges is that since most training data is located in a small subspace of the high dimensional data space, training the score network in low data density regions can be challenging.
Therefore, Song and Ermon (2019) proposed to perturb the data space using a scheduled Gaussian noise and learn the score function using the perturbed data following Vincent (2011). This approach is also called “denoising score matching,” and as shown in Vincent (2011), if the perturbation is small enough, learning the score function of a perturbed data distribution ($q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)$) is almost surely equivalent to learning the actual score function of $p(\mathbf{x})$, i.e., $\nabla_{\mathbf{x}}\log q_{\sigma}(\tilde{\mathbf{x}}|\mathbf{x})\approx\nabla_{\mathbf{x}}$. Therefore, the updated learning objective is as follows:
$\begin{split}&\mathbbm{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\mathbbm{E}_{\tilde{\mathbf{x}}\sim q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)}\left\lVert\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)-s_{\theta}(\tilde{\mathbf{x}})\right\rVert^{2}_{2}\right]\\ =&\mathbbm{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\mathbbm{E}_{\tilde{\mathbf{x}}\sim\mathcal{N}\left(\mathbf{x},\sigma^{2}\mathbbm{I}\right)}\left\lVert\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}-s_{\theta}(\tilde{\mathbf{x}})\right\rVert^{2}_{2}\right],\\ \end{split}$ (2.28)
where, $\tilde{\mathbf{x}}$ is the perturbed data with Gaussian distribution with standard deviation $\sigma$. Song and Ermon (2019) further proposed to use a variance scheduling for data perturbation. Formally, we can define a set of standard deviation of data perturbation, $\{\sigma_{t}\}^{L}_{t=1}$, that satisfies $\frac{\sigma_{1}}{\sigma_{2}}=\cdots=\frac{\sigma_{L-1}}{\sigma_{L}}>1$. Specifically, we start from a large enough data perturbation that can cover (or generate data in) the low data density regions in the data space, and then we gradually decrease the perturbation so that the learned score function can match with the actual score function. Intuitively, we can interpret the large perturbation as learning the general trend or landscape of the score function. In contrast, we can interpret the small perturbation as learning the details of the actual score function.
Then, the score network is trained given the current noise level ($\sigma_{t}$) and this is called the “Noise Conditional Score Network.” For a given $\sigma$, the updated learning objective is:
$\begin{split}\mathbbm{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\mathbbm{E}_{\tilde{\mathbf{x}}\sim\mathcal{N}\left(\mathbf{x},\sigma^{2}\mathbbm{I}\right)}\left\lVert\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}-s_{\theta}(\tilde{\mathbf{x}},\sigma)\right\rVert^{2}_{2}\right].\\ \end{split}$ (2.29)
Once the score network is trained, data generation can be accomplished through an iterative method known as Langevin dynamics, or Langevin Monte Carlo (Parisi, 1981; Grenander and Miller, 1994). Langevin dynamics is originally formulated to describe the behavior of molecular systems, such as Brownian motion. A more detailed derivation of equations for Langevin dynamics in the generative model setting is presented in Appendix B. Given a fixed step size $\epsilon>0$, and an initial random noise $\mathbf{x}$, the following iterative procedure can be used:
$\mathbf{x}_{i+1}\leftarrow\mathbf{x}_{i}+\nabla_{x}\log p(\mathbf{x})\cdot\epsilon+\sqrt{2\epsilon}\cdot\mathbf{z}_{i},$ (2.30)
where $\mathbf{z}_{i}\sim\mathcal{N}(0,I)$. Since the values from the trained score network should match the true score, we can use the following iterative procedure to draw samples from the distribution of interest:
$\mathbf{x}_{i+1}\leftarrow\mathbf{x}_{i}+s_{\theta}(\mathbf{x}_{i})\cdot\epsilon+\sqrt{2\epsilon}\cdot\mathbf{z}_{i}.$ (2.31)
2.7 Diffusion Models
Diffusion models are a class of DGMs that have shown a remarkable ability to generate high-quality samples across various domains. The core concept behind diffusion models is inspired by the physical process of “diffusion,” where particles move from areas of higher concentration to lower concentration until they reach an equilibrium. Diffusion models work by gradually adding noise to data until it becomes indistinguishable random noise (known as a forward diffusion process, or diffusion process), and then learning the reverse of this process of denoising the data (known as a reverse diffusion process, or generation process).
Figure 5: Schematic overview of diffusion model
Figure 5 shows the general framework of the diffusion model. We start from $\mathbf{x}_{0}$ which is the real data. At each diffusion step, we add a small noise. After a sufficiently large number of steps (here we denote it as $k$), the data is nearly complete noise as shown in $\mathbf{x}_{k}=\mathbf{z}$. The forward process of adding noise is given, and the main ‘learning’ part of the diffusion model is to learn the reverse process which transforms the complete noise $\mathbf{z}$ to the data $\mathbf{x}_{0}$. This is similar to VAE, GAN, and Normalizing Flows that we start from a complete noise (or well-known distribution such as Gaussian distribution or uniform distribution) and we learn to transform the noise into the data distribution $p(\mathbf{x})$. The main difference is the idea inspired by the physical “diffusion” process and denoising process which recovers the data from a complete noise.
Ho et al. (2020) presents the Denoising Diffusion Probabilistic Model, or DDPM, which extends the diffusion model framework by incorporating a denoising autoencoder-like architecture. This architecture allows for more efficient learning of the reverse diffusion process, enabling the model to generate high-fidelity samples from complex data distributions.
The forward diffusion process is designed to corrupt the original data $\mathbf{x}_{0}$ over a series of time steps $t=1,2,\ldots,T$. At each step $t$, Gaussian noise is added to the data, which is mathematically described by the conditional distribution $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. This distribution specifies how the data at step $t$ is generated from the data at the previous step $t-1$.
Formally, the forward process is defined as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbbm{I}),$ (2.32)
where $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ follows the conditional Gaussian distribution with mean $\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}$ and covariance $\beta_{t}\mathbbm{I}$, $t$ is the diffusion timestep, and $\beta_{t}$ is a time-dependent coefficient that controls the amount of data retained from the previous step where $0<\beta_{t}\ll 1$. The choice of $\sqrt{1-\beta_{t}}$ and $\beta_{t}$ ensures that the variance of each diffusion step is maintained at 1. This is important because it stabilizes the diffusion process, preventing the variance from either exploding or vanishing over time.
Based on the reparameterization trick, the transition from $\mathbf{x}_{t-1}$ to $\mathbf{x}_{t}$ can be also expressed as:
$\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\beta_{t}\boldsymbol{\epsilon}_{t},$ (2.33)
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,I)$ is Gaussian noise with mean 0 and identity covariance matrix.
From Equation (2.32), we can derive the equation for $q(\mathbf{x}_{0:T})$ representing the diffusion trajectory from the original data $\mathbf{x}_{0}$ to the complete noise $\mathbf{x}_{T}=z$ based on properties of Markov chain and chain rules as follows:
$\begin{split}&q\left(\mathbf{x}_{0:T})=q(\mathbf{x}_{0}\right)\prod_{t=1}^{T}q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right),\\ &\text{or,}\\ &q\left(\mathbf{x}_{1:T}|\mathbf{x}_{0}\right)=\prod_{t=1}^{T}q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right).\\ \end{split}$ (2.34)
The reverse diffusion process (the denoising process) represents the generative distribution parameterized with $\theta$. The joint distribution $p_{\theta}\left(\mathbf{x}_{0:T}\right)$ is defined as a Markov chain with Gaussian transition distributions, $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$ starting at $p\left(\mathbf{x}_{T}\right)=\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbbm{I})$ as follows:
$p_{\theta}\left(\mathbf{x}_{0:T}\right)=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right),$ (2.35)
where
$p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t},t)\right).$ (2.36)
Similar to VAE and Normalizing Flows (models using explicit densities for training), the training objective is to maximize the likelihood or minimize the negative log-likelihood
$\theta^{*}=\arg\min_{\theta}\mathbbm{E}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right].$ (2.37)
Instead of directly training the model with negative log-likelihood, we can consider $q$ as the approximate posterior and use the variational bound on negative log-likelihood to train the model. A detailed derivation is presented in Appendix A and Ho et al. (2020). As a result, we can find an upper bound of the optimization
$\begin{split}\mathbbm{E}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right]\leq\mathbbm{E}\left[\underbrace{D_{KL}\left(q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)||p\left(\mathbf{x}_{T}\right)\right)}_{L_{T}}+\sum_{t=2}^{T}\underbrace{D_{KL}\left(q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)||p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)\right)}_{L_{t-1}}\underbrace{-\log p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}_{L_{0}}\right].\end{split}$ (2.38)
Therefore, the overall loss function of minimizing the negative log-likelihood in Equation (2.37) is decomposed into several losses, $L_{T}$, $L_{t-1}$, and $L_{0}$. Here, $L_{T}$ is constant since both $q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)$ and $p\left(\mathbf{x}_{T}\right)$ are fixed, and therefore, we can ignore this term. Also, in Ho and Ermon (2016), $L_{0}$ is explicitly defined by using the characteristics of the image generation problem, and as a result, $L_{0}$ can be interpreted as a reconstruction loss of a problem-specific decoder. As a result, the actual learning process of the diffusion model is related to $L_{t-1}$.
$L_{t-1}$ measures the KL-divergence of $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$ from $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$. The diffusion process, $q$, represents the process of adding small noise to the data; i.e., given a less noisy data $\mathbf{x}_{t-1}$, the distribution of a more noisy data $\mathbf{x}_{t}$. The first term, $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$, represents the true denoising process which is derived from the definition of $q$ given the true data without noise, $\mathbf{x}_{0}$. What the diffusion models try to learn is the denoising process $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$; i.e., given a more noisy data $\mathbf{x}_{t}$, the distribution of a less noisy data $\mathbf{x}_{t-1}$. As a result, $L_{t-1}$ captures the distributional difference between the true denoising process (given the true data) and the approximated denoising process (without the true data).
Since the diffusion process follows Gaussian distribution, the true reverse process, $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$, can be assumed to follow a Gaussian distribution if $T$ is sufficiently large, or $T\rightarrow\infty$. Let $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)=\mathcal{N}(\mathbf{x}_{t-1};\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0}),\tilde{\beta}_{t}\mathbbm{I})$. To derive explicit form of $\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0})$ and $\tilde{\beta}_{t}$, first, we should derive a closed-form equation for sampling $\mathbf{x}_{t}$ at an arbitrary timestep $t$ from Equation (2.32) as follows:
Then, Ho et al. (2020) found that using $L_{simple}$ results in better training in empirical testing.
$L_{simple}=\mathbbm{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\lVert\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}}_{t}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},t\right)\right\rVert_{2}^{2}\right].$ (2.39)
2.8 Discussion
Different classes of Deep Generative Models (DGMs) use various approaches to model data distributions. Since all models have their imperfections, it’s crucial to understand the commonly known pros and cons of each class. Typically, we expect three key things from DGMs: 1) High-Quality Sample Generation — The ability to produce samples that are indistinguishable from real data; 2) Mode Coverage and Sample Diversity — The capacity to capture all the variations in the data, including minority modes, ensuring that the generated samples are diverse and representative of the entire data distribution; and 3) Fast and Computationally Efficient Sampling— The ability to generate samples quickly without requiring extensive computational resources.
However, most DGMs cannot yet simultaneously satisfy these three key requirements, a challenge often referred to as the Generative Learning Trilemma (Xiao et al., 2021). This trilemma highlights the inherent trade-offs between high-quality sample generation, mode coverage and diversity, and fast, computationally efficient sampling. For example, Generative Adversarial Networks (GANs) have the advantage of generating high-quality samples rapidly. As discussed in Section 2.4, GANs are renowned for producing realistic samples through a single forward pass of the generator network, making them computationally efficient at inference time. However, common shortcomings of GANs include mode collapse, where the model fails to capture the full diversity of the data distribution, generating samples from only some modes but not all. On the other hand, Variational Autoencoders (VAEs) and Normalizing Flows usually perform well in terms of mode coverage and fast computation time. VAEs promote diverse and representative sampling by modeling the entire data distribution, and they allow for quick sample generation by direct decoding from latent variables. Normalizing Flows also enables efficient sampling and exact density estimation. However, the quality of the samples generated by VAEs and Normalizing Flows may not be as high as those produced by other models. Diffusion Models present another approach, capable of covering diverse modes and generating high-quality samples. Recent advances have enabled diffusion models to produce samples that rival or even surpass GANs in terms of quality and diversity. However, diffusion models typically consist of a large number of denoising steps (often ranging from 500 to 1000), which require substantial computation to generate a single sample. This makes the sampling process slow and computationally expensive compared to other models.
As a result, understanding these trade-offs of using each class of DGM is essential when applying DGMs in practice. Since no current model perfectly satisfies all three key requirements, researchers should choose the model class that best aligns with the specific needs of their application, whether that be sample quality, diversity, or computational efficiency. Additionally, some previous research has explored combining different modeling approaches from several DGMs, such as Xiao et al. (2021) which combined Diffusion models with GAN, Grover et al. (2018) which combined Normalizing Flows with GAN, and Zhang and Chen (2021) which combined Normalizing Flows with Diffusion model. These examples demonstrate that by integrating different DGMs, researchers can also address the limitations imposed by using a single DGM.
Deep Generative Models in Different Areas of Transportation Research
DGM Application in Transportation Research Data Generation (Section 3.1) Estimation and Prediction (Section 3.2) Unsupervised Representation Learning (Section 3.3) Scenario and Synthetic Data Generation (Section 3.1.1) Trajectory Generation (Section 3.1.2) Missing Data Imputation (Section 3.1.3) Agent-level Analysis (Section 3.2.1) Link-level Analysis (Section 3.2.2) Region-level Analysis (Section 3.2.3) Feature Extraction and Classification
Figure 6: An overview of DGMs in transportation research.
DGMs have become important in current research for their ability to model complex data distributions and generate data samples that closely mimic real-world observations. Particularly, in transportation research, DGMs offer tools to simulate, predict, and optimize transportation systems effectively (Huynh and Phung, 2021; Lin et al., 2023a; Yan and Li, 2023). Therefore, in this section, we provide a systematic review of the applications of DGMs in transportation research.
This review focuses on three main areas where DGMs have significantly impacted transportation research: data generation, estimation and prediction, and unsupervised representation learning. By examining their applications in these areas, we gain a clear understanding of how DGMs impacted and will impact transportation research. An overview of this section is shown in Figure 6.
  1. 1.
    DGM for data generation: The core functionality of DGMs lies in their ability to generate realistic synthetic data that closely replicate real-world conditions. This capability is particularly valuable in transportation research, where collecting large, high-quality datasets can be challenging and costly. Through the generation of diverse traffic scenarios, DGMs enable researchers to investigate situations that are difficult to capture through traditional data collection methods.
  2. 2.
    DGM for estimation and prediction: Beyond generating data, DGMs are also highly effective for estimation and prediction tasks. DGMs learn the underlying distribution of the given (training) data, and as a result, they can model the inherent uncertainties and variations in traffic data. This capability allows DGMs to provide probabilistic estimations and forecasts of traffic states. The probabilistic approach enhances the accuracy, robustness, and reliability of predictions, which support better decision-making in transportation planning and operation.
  3. 3.
    DGM for unsupervised representation learning: Most DGMs are designed to learn the transformation function from a sample from a simple distribution (also known as a base distribution) to a target distribution. The ‘noisy’ data sample from the base distribution can be considered as a latent representation of the synthetic data sample. These latent representations can be used for various analyses, such as classifying transportation modes, detecting traffic anomalies, or understanding driving behaviors. This ability to derive insights without requiring labeled data makes DGMs powerful tools in unsupervised learning scenarios.
3.1 DGM for Data Generation
The rapid urbanization of society requires effective traffic management systems to reduce congestion, improve urban planning, and enhance transportation network efficiency (Xu et al., 2015). These systems are fundamentally built on large volumes of high-quality traffic data. Sufficient data volume is essential for accurate analyses and modeling of complex traffic behaviors and patterns, while data quality ensures precise and reliable insights and applications. Despite advancements in technology that have improved data collection capabilities in transportation, traffic sensors often face issues of sparsity and unreliability, making it challenging to obtain adequate, high-quality data in real-world scenarios (Chen et al., 2003; Ni and Leonard, 2005). DGMs address these challenges by both increasing the volume and improving the quality of traffic datasets. They can generate synthetic data to supplement existing datasets, effectively expanding their size. Additionally, DGMs replicate real-world conditions with high fidelity, which enhances the dataset’s qualitative aspects. These capabilities are particularly valuable for transportation research, where traditional data collection methods may be costly, time-consuming, and impractical, especially for capturing rare or non-reproducible scenarios. For example, synthetic data generated by DGMs can simulate a range of traffic conditions, from typical daily commutes to rare events like accidents or extreme weather. Consequently, the synthetic data from DGMs is crucial for developing robust traffic models and systems that can handle a wide range of scenarios. In the following sections, we will explore data generation through DGMs in three key areas: scenario and synthetic data generation, trajectory generation, and missing data imputation.
3.1.1   Scenario and Synthetic Data Generation
Scenario and synthetic data generation involve using DGMs to produce a wide range of synthetic data. This process includes creating diverse driving scenarios to test and evaluate autonomous vehicles, synthesizing demographic and travel behavior data to represent different populations in studies, and generating synthetic traffic data for training other machine learning models for traffic operation and management. Additionally, DGMs can generate rare or unexpected traffic situations to improve anomaly detection and response strategies. These capabilities allow researchers to simulate various conditions and events which can help enhance the robustness and accuracy of analysis. This section highlights how the DGMs enable comprehensive testing, analysis, and improvement of transportation networks through effective scenario and synthetic data generation.
3.1.1.1   Driving Scenario Generation
Many studies in this section have focused on driving scenario generation, which involves creating diverse and realistic driving situations to test and evaluate autonomous vehicles and traffic management systems (Ghosh et al., 2016; Yun et al., 2019; Singh et al., 2023; Ye et al., 2022; Jin et al., 2023; Dong et al., 2023; Xu et al., 2023b; Huang et al., 2024; Li et al., 2024d). Most of these studies utilize additional information as conditions for their generative models, such as car-following theory, safety critics, or physical restrictions, to generate more realistic and high-performing models. For example, Chen et al. (2022a) developed a two-step hybrid driving model that combines model-based controllers with Generative Adversarial Imitation Learning (GAIL) for traffic simulation. The model generates high-level driver traits used as parameters for low-level model-based controllers to simulate human-like driving in multi-agent traffic scenarios. Zhong et al. (2023) presented a framework using a conditional diffusion model to generate realistic and controllable traffic simulations. The key benefit of the model is that it allows users to specify trajectory properties, such as reaching a goal or following speed limits while ensuring that the generated trajectories are physically feasible and realistic through the guidance of signal temporal logic rules. Similarly, Sun et al. (2023) introduced a data-driven method DriverSceneGen for generating diverse and realistic driving scenarios to address data limitations. DriverSceneGen employed a diffusion model to create initial driving scenes in a rasterized Bird-Eye-View (BEV) format and used a simulation network to predict multiple future scenarios based on the initial setup. Furthermore, Xu et al. (2023a) developed DiffScene, a diffusion-based framework designed to generate safety-critical driving scenarios for evaluating and enhancing autonomous vehicle safety. DiffScene used a diffusion model to approximate the distribution of low-density areas in traffic data, creating realistic safety-critical situations. These scenarios were further optimized using guided adversarial objectives to maintain both realism and safety-criticality. Additionally, Zhang et al. (2020a) employed the conditional GAIL (cGAIL) framework to study taxi drivers’ decision-making strategies. The method transferred knowledge across different drivers and locations to improve policy accuracy, improving taxi drivers’ income and service quality. These studies have demonstrated the capability of using DGMs in driving scenario generation by providing more diverse and realistic scenarios for traffic management and Autonomous Vehicle development.
3.1.1.2   Population Synthesis
Population synthesis is a critical component of transportation research that involves generating synthetic demographic and travel behavior data to accurately represent populations. This synthetic data is essential for effectively modeling and analyzing transportation systems. Many studies have utilized VAEs or GAN-based models to solve the problems (Borysov et al., 2018; Badu-Marfo et al., 2020; Mensah et al., 2022; Johnsen et al., 2022; Badu-Marfo et al., 2022). For instance, Borysov et al. (2019) presented a VAE-based approach for population synthesis to address common challenges, such as scalability and the issue of sampling zeros in previous methods. The proposed method fits the full joint distribution of high-dimensional data, enabling the generation of more comprehensive synthetic datasets. Validation with a large-scale Danish trip diary dataset demonstrated the method’s effectiveness in enhancing the realism and diversity of synthetic populations. Kim and Bansal (2023) also tackled the challenges of structural zeros and the generation of feasible, diverse synthetic populations. They employed customized loss functions for the VAE and Wasserstein GAN (WGANs) and validated their approach using a large household travel survey dataset from South Korea. Similarly, Garrido et al. (2020) proposed a VAE and WGAN-based method to predict rare feature combinations in population synthesis. These models generate synthetic populations that include rare and diverse feature combinations throughout capturing the joint distribution of high-dimensional attributes. Validation with the Danish National Travel Survey data demonstrated notable improvements in generating realistic and varied synthetic populations.These studies collectively highlight the versatility and robustness of DGMs in population synthesis. They underscore the importance of DGMs in creating realistic and diverse synthetic data, which is crucial for accurate transportation modeling and analysis.
3.1.1.3   Synthetic Data and Model Training
The DGM can also be used to generate various types of synthetic traffic data for traffic modeling and training purposes. These synthetic datasets enable researchers to simulate traffic scenarios and train models without relying solely on real-world data, which can be expensive and difficult to collect. In related research, instead of using basic models, most researchers combine various deep learning architectures or incorporate additional information to further improve performance (Chen et al., 2019; Ozturk et al., 2020; Wu et al., 2020; Chen et al., 2021c; Zhou et al., 2023; Rong et al., 2023b; Kumar et al., 2023; Devadhas Sujakumari and Dassan, 2023). For traffic data generation, Li et al. (2022a) proposed the Attentive Dual-Head Spatial-Temporal GAN (ADST-GAN) to generate crowd flows. The model integrates attentive temporal and spatial mechanisms by combining ConvLSTM, self-attention, and a dual-head discriminator within the GAN framework. These components capture complex spatial-temporal dependencies in crowd flow data and mitigate overfitting, ensuring that the synthetic data generated closely mimics real-world traffic patterns. To address the complexity of modeling Origin-Destination (OD) networks, Rong et al. (2023a) introduced a large-scale OD generation method using a graph-denoising diffusion model. This approach simplifies the generation process by decomposing it into two stages: first, generating the network topology, and second, generating the edge weights. By tackling these stages separately, the model accurately generates realistic OD data, making the analysis of large-scale OD networks more manageable. For model training, Dewi et al. (2022) applied a Deep Convolutional GAN (DCGAN) to generate synthetic traffic sign images. This approach captures the distribution of real traffic sign images and produces new samples to augment the original dataset. The researcher further combined these synthetic images with real ones to improve the performance of deep recognition models, such as CNN and ResNet 50. Additionally, Zhu et al. (2022) employed the Spatio-Temporal Dependencies GAN (STDGAN) to generate high-quality synthetic traffic volume data for prediction tasks. Specifically, STDGAN captures implicit variation patterns in traffic volume based on information from traffic speed and occupancy data. Experiment results indicate that enriching the training dataset with the synthetic volume data led to more robust and accurate performance in prediction models. These examples demonstrate the ability of DGMs to generate various traffic data that support traffic management and enhance model training, offering valuable resources for transportation research and applications.
3.1.1.4   Anomaly Data Generation
Anomaly data generation is crucial for creating rare or unexpected traffic situations, which are essential for testing and evaluating systems under extreme conditions. Generating such data helps develop robust models that can handle these rare events effectively. In this research field, GAN-based models are widely adopted due to their flexible and high performance in generating synthetic data for anomaly situations (Islam et al., 2021; Zarei and Hellinga, 2021; Li et al., 2024a; Chen et al., 2024). For instance, Cai et al. (2020) utilized a DCGAN to generate synthetic crash data from learning the distribution of crash-related traffic data. The synthetic data was then combined with real data to create a balanced dataset, which was used to train various crash prediction algorithms, including Logistic Regression, SVM, ANN, and CNN. Similarly, Chen et al. (2022d) employed two GAN variants—Conditional GAN (CGAN) and Gaussian Discriminant Analysis GAN (GDAGAN)—to tackle the challenge of imbalanced traffic collision datasets. These generative models generated additional samples for underrepresented classes, thereby balancing the dataset and enhancing the performance of classifiers trained on them. Huo et al. (2021) introduced the Text-to-Traffic GAN (T2GAN), which generates realistic traffic situations by integrating traffic data with semantic information from social media. In the proposed model, CGAN was used to generate traffic scenarios based on text inputs from social media, which enables the creation of varied and contextually rich traffic situations. To tackle data imbalance in traffic crash severity modeling, Chen et al. (2024) developed a method based on Conditional Tabular GAN (CTGAN) combined with Random Under-sampling (CTGAN-RU). This approach generated synthetic crash data while accounting for both continuous and discrete characteristics of the imbalanced dataset. The integration of various conditions helped create a more balanced training set and improving the accuracy and reliability of severity prediction models. These studies demonstrate the importance of using DGMs to generate anomaly data, which helps address the imbalanced data problem and improves the robustness of traffic management and safety systems.
3.1.2   Trajectory Generation
Trajectory generation creates realistic movement patterns for vehicles and pedestrians, which is essential for developing accurate traffic simulations and urban planning tools. By leveraging DGMs, researchers can produce synthetic trajectories that mimic real-world behaviors of traffic agents, enabling more robust analysis and testing of transportation systems. These generated trajectories help understand traffic dynamics, evaluate transportation policies, and enhance infrastructure design. These trajectories can be further classified based on their scale and scope, such as micro and macro scales. Micro-scale trajectories focus on detailed, short-term movement patterns of individual agents in localized areas, like intersections, while macro-scale trajectories represent broader, long-term traffic flow across extensive road networks or regions. The classification into micro and macro scales significantly impacts their use cases, such as collision avoidance or urban planning, and determines the complexity of the problems they address.
3.1.2.1   Micro-scale Trajectory Generation
Micro-scale trajectory generation focuses on producing detailed, short-term predictions for localized areas, such as intersections or specific road segments. These trajectories are critical for applications requiring precise and immediate responses, like collision avoidance, maneuver planning in dense urban settings, and real-time navigation adjustments. Numerous studies have leveraged various DGMS to produce realistic micro-scale trajectories (Krajewski et al., 2018; Ding et al., 2019; Krajewski et al., 2019; Gong et al., 2023; Yang et al., 2024). For example, Agarwal et al. (2020) used normalizing flows in generating trajectories for Autonomous Vehicles (AVs) through the FlowPlan method. This approach employs a Conditional VAE (CVAE) to encode human expert trajectories into a low-dimensional latent space. A Neural Autoregressive flow (NAF) then maps this latent space to a Boltzmann distribution over the AV control space, effectively capturing complex, multimodal distributions of feasible control trajectories conditioned on the scene context. Rempe et al. (2023) used a guided diffusion model to generate realistic pedestrian trajectories. The proposed model integrates with a physics-based humanoid controller, forming a closed-loop system capable of simulating large crowd movements across various scenarios. By combining the guided diffusion model with physical simulation, it accurately generates pedestrian trajectories that reflect real-world behavior. Similarly, Niedoba et al. (2024) employed a diffusion-based generative model to simulate realistic traffic scenarios involving multiple interacting agents. To be more specific, it diffuses the trajectories of all agents in a scene jointly, based on past, present, or future state observations. This method allows for the creation of diverse and realistic traffic scenarios, including rare, safety-critical events. These studies highlight how DGMs can produce highly detailed and realistic micro-scale trajectories, which are essential for applications that require precise and immediate response capabilities in traffic systems.
3.1.2.2   Macro-scale Trajectory Generation
Macro-scale trajectory generation covers broad spatial and temporal scopes, including extensive areas such as entire road networks or large regions over longer periods. The trajectories are crucial for strategic planning, traffic management, and long-term behavioral analysis. By examining origin-destination pairs, travel time distributions, and overall traffic flow patterns, macro-scale trajectories provide valuable insights into larger trends and systemic issues. The information is vital for urban planning, optimizing public transportation systems, and enhancing the efficiency of traffic networks. Numerous studies have focused on generating macro-scale trajectories with conditional information based on the enhanced DGM by various deep-learning models. (Chen et al., 2021c, b; Zhang et al., 2022b; Zhu et al., 2024a; Wei et al., 2024; Wang and Kankanhalli, 2024; Zhu et al., 2024b). For example, Xiong et al. (2023) introduced TrajSGAN, a model designed to generate human mobility trajectories for urban behavior analysis and decision-making. It integrates an attention-based generator for trajectory location prediction with a rollout module and a CNN-based discriminator to assess the spatial structure of the generated trajectories. This integration ensures that the results are realistic and semantically guided. TrajSGAN effectively reduces divergence in spatial-related metrics and can be utilized to study epidemic diffusion with high accuracy. Similarly, Choi et al. (2021a) developed TrajGAIL, a GAIL framework for generating urban vehicle trajectories. The model combines the capabilities of GAIL with RNNs to learn the underlying distributions of urban vehicle trajectory data. TrajGAIL allows for the generation of synthetic vehicle trajectories that closely resemble real-world patterns, even from limited observations. Sun and Kim (2023) introduced MAGAIL-VL, a data-driven, network-wide traffic simulation framework considering both Vehicle and Link agents. Specifically, MAGAIL-VL learns the distribution of vehicle movements and link states from observed data, enabling the generation of realistic traffic scenarios across an entire network. Recently, diffusion-based models have become popular for generating macro-scale trajectories (Zhu et al., 2024a; Wei et al., 2024; Wang and Kankanhalli, 2024; Zhu et al., 2024b). For example, Zhu et al. (2024a) introduced a novel approach DiffTraj, which generates high-quality synthetic GPS trajectories using a spatial-temporal diffusion probabilistic model. DiffTraj model combines residual blocks and multi-scale feature fusion in U-Net architecture to accurately predict noise levels during the reverse denoising process, enhancing its performance in capturing complex spatial-temporal dependencies and generating realistic synthetic trajectories. Similarly, Wei et al. (2024) presented Diff-RNTraj, a diffusion-based model designed to generate trajectories that are both geographically accurate and adhere to the logical constraints of the road network. The model combines a continuous diffusion framework with a pre-training strategy to manage the hybrid nature of road network-constrained trajectory data, resulting in realistic and usable synthetic trajectories. These studies highlight the diverse applications of DGMs in producing macro-scale trajectories. They provide essential tools and insights for improving transportation systems through strategic planning and management.
3.1.3   Missing Data Imputation
The missing data imputation in transportation aims to reconstruct original data $X$ from missing or corrupted data $\hat{X}$ with conditional information $\phi$ such as known historical data, geometry configuration, and external factors ($\hat{X}=f(X,\phi)$). This process inherently involves data generation, as it creates new values to fill gaps in existing datasets, thereby enhancing the volume and quality of data available for analysis. In traditional imputation methods, missing data is often challenging to reconstruct due to the complex spatio-temporal dependencies in traffic data, which limit the accuracy and robustness of simple statistical methods. In recent years, DGMs have become powerful tools for handling the complexities of traffic data imputation. Their ability to model the underlying distribution of traffic data enables them to generate realistic and accurate reconstructions of missing entries across various scenarios. DGMs can also integrate a wide range of conditional information, including the spatio-temporal dependencies of traffic data, to enhance imputation performance (Zhang et al., 2024a; Cai et al., 2023; Yang et al., 2021; Zhang et al., 2021c; Peng et al., 2023; Shin et al., 2023; Duan et al., 2024; Chen et al., 2022b). Traditionally, missing data has been categorized into three types: Missing Completely at Random (MCAR), Missing at Random (MAR), and Missing Not at Random (MNAR) (Lin and Tsai, 2020; Hasan et al., 2021). However, the unique spatio-temporal characteristics of traffic data require a reevaluation of these categories. To address this, Chan et al. (2023) proposed a new classification specifically for transportation research, dividing missing data into Fiber Missing Data, Block Missing Data, and Random Missing Data. In this paper, we adopt this updated categorization to better address the specific challenges of traffic data imputation.
3.1.3.1   Fiber Missing Data
Fiber Missing Data occurs when there is a sudden, temporary failure of data-acquisition devices, leading to gaps in data collection that can be short-term or long-term. Many advanced models have been developed to address this issue effectively (Li et al., 2018; Zhang et al., 2021c, b; Huang and Chen, 2022; Yuan et al., 2022; Shen et al., 2022; Cai et al., 2023; Zhang et al., 2023b; Hou et al., 2023; Cai et al., 2023; Li et al., 2024c; Duan et al., 2024; Zhang et al., 2024a). For instance, Han et al. (2020) developed the Content-Aware GAN (CA-GAN) to handle missing traffic data in time series. In this method, traffic data were considered as tensors to capture their spatial-temporal correlations. The CA-GAN learns to generate realistic traffic patterns, effectively completing gaps in speed data. Similarly, Xu et al. (2021) introduced the Graph Aggregate GAN (GA-GAN), which combines the strengths of Graph Sample and Aggregate (GraphSAGE) with a GAN to impute missing traffic speed data. GraphSAGE aggregates information from neighboring nodes in the road network to capture spatial-temporal correlations, which are then used by a Wasserstein GAN (WGAN) to generate the missing data. The GA-GAN has been tested in both fiber and random missing data environments, demonstrating its versatility. Zhang et al. (2021b) proposed the Gated GAN (GaGAN) to address missing speed data on signalized roads. GaGAN incorporates attention mechanisms into graph convolutional operations and GRUs to learn spatio-temporal correlations within each signal cycle. This approach is effective for both fiber and random missing data scenarios. In another study, Liu et al. (2023) introduced PriSTI, a conditional diffusion framework for spatio-temporal data imputation. Specifically, the PriSTI uses a diffusion-based noise estimation module and a spatio-temporal feature extraction process to enhance its performance under high missing rates and diverse spatial configurations. It has been evaluated in environments with both fiber and random missing data, showing significant robustness. These models demonstrate the capabilities of DGMs in addressing the challenges posed by fiber missing data, ensuring the continuity and quality of traffic datasets even in the face of temporary data collection failures.
3.1.3.2   Block Missing Data
Block Missing Data occurs when there are no data-acquisition detectors in the area of interest, leading to complete loss in the dataset for the region. This type of missing data is particularly challenging to address because of the absence of any observations over an extended spatial and temporal range. Advanced DGMs have been developed to tackle this challenge, though most studies do not focus exclusively on block missing data. Instead, they often evaluate model performance under various conditions, including fiber and random missing data scenarios(Li et al., 2018; Zhang et al., 2021c; Yuan et al., 2022; Hou et al., 2023; Li et al., 2023b; Zhang et al., 2024a). For example, Yuan et al. (2022) enhanced the performance of GAN in traffic flow imputation by designing generative and center losses. This approach was evaluated across all three types of missing data environments: Fiber, Block, and Random Missing Data. The proposed method demonstrated significant improvements in imputing missing traffic data under these conditions. In another study, Li et al. (2018) combined a 3D Convolutional Neural Network with GAN (3wwDConvGAN) to address missing traffic data. Unlike traditional methods that may not fully exploit the spatial-temporal features of historical data, 3DConvGAN uses a fractional strided 3D CNN in both the generator and discriminator to enhance imputation performance. The model’s effectiveness was also tested in environments with Fiber, Block, and Random Missing Data, highlighting its versatility. Zhang et al. (2021c) introduced SA-GAIN, a Self-Attention Generative Adversarial Imputation Network, designed to impute missing traffic flow data. By incorporating a self-attention mechanism in GAN, SA-GAIN effectively captures important features across spatial and temporal dimensions. The model was tested in both fiber and block missing data scenarios, demonstrating robust performance. Hou et al. (2023) developed MissII, a model that combines spatial interaction models with GANs for traffic data imputation. This method first estimates traffic flow using theory-guided models that leverage environmental and social factors to capture complex traffic dynamics. The estimated traffic flow data are then used as real samples to guide the GAN’s training process to improve the imputation accuracy. These studies illustrate the effectiveness of DGMs in addressing block missing data, which presents unique challenges due to the complete absence of data in specific areas. By leveraging spatio-temporal modeling and integrating advanced neural network architectures, these models provide comprehensive solutions to ensure continuity and quality in traffic data even when entire sections of data are missing.
3.1.3.3   Random Missing Data
Random Missing Data occurs when the data-acquisition detectors fail unpredictably in the spatio-temporal domain, which results in gaps in data with little to no correlation among the missing entries. Due to its unpredictable nature, this scenario presents significant challenges and is commonly used to evaluate the performance of imputation models (Li et al., 2018; Chen et al., 2021a, 2022b; Wu et al., 2022a; Qin et al., 2021; Kazemi and Meidani, 2021; Tu et al., 2021; Zhang et al., 2021b; Xu et al., 2021; Shen et al., 2022; Yuan et al., 2022; Yang et al., 2022; Zhang et al., 2022a; Wang et al., 2023; Liu et al., 2023; Li et al., 2023b; Huang et al., 2023; Zhang et al., 2023b; Cai et al., 2023; Hou et al., 2023; Cai et al., 2023; Li et al., 2024c; Zheng et al., 2024; Duan et al., 2024; Zhang et al., 2024b, a). For example, Chen et al. (2019) developed a GAN-based method to impute missing traffic flow data by combining real and synthetic data. This approach utilizes a parallel data paradigm where the GAN generates synthetic data that are used alongside real data to train the imputation model, enhancing its ability to fill in gaps accurately. Boquet et al. (2019, 2020) used a VAE-based method for imputing missing traffic data, aimed at improving the accuracy of traffic forecasting systems impacted by sensor or system failures. The unsupervised approach learns the underlying data distribution from a latent space, resulting in improved post-imputation performance and effective data augmentation. To fully exploit the benefits of both VAE and GAN in data imputation, Shin et al. (2023) presented an Adversarial Autoencoder (AAE)-based model that leverages spatio-temporal feature extraction to address missing traffic data. AAE combines the principles of VAE and GAN to capture the complex dependencies within traffic data, providing robust imputation capabilities even in scenarios with high rates of missing data. Yang et al. (2021) introduced the Spatio-Temporal Learnable Bidirectional Attention GAN (ST-LBAGAN) for imputing missing traffic data. This model integrates U-Net and bidirectional attention mechanisms into a GAN-based framework to capture spatio-temporal dependencies for improving imputation performance. Additionally, Li et al. (2023a) proposed the Time Series Generative Adversarial Imputation Network (TGAIN) to tackle the challenge of imputing time series data. This model is designed to learn the multi-state distribution of missing time series data under conditional vector constraints. By employing a multiple imputation strategy, it effectively handles the uncertainty inherent in the imputation process. Recently, Zheng et al. (2024) used traffic domain knowledge and a denoising diffusion model to develop the Doubly Physics-Regularized Denoising Diffusion Model (DPRDDM) for recovering corrupted traffic speed data. This model demonstrates robustness in handling various types of noise, including Gaussian white noise, random corrupted noise, spatially correlated noise, temporally correlated noise, and a mixture of these noise types. These models illustrate the diverse applications and robust capabilities of DGMs in addressing the challenges posed by random missing data. By effectively capturing and utilizing spatio-temporal dependencies, these models ensure accurate and reliable data imputation even in the most unpredictable scenarios.
3.2 DGM for Estimation and Prediction
One of the significant aspects of DGMs in transportation studies is their capability to learn and accurately model the distribution of input data. This foundational property enables DGMs to handle complex datasets and is critical for advanced analytical tasks such as traffic data generation, prediction, and classification. Another critical yet less explored advantage of DGMs is their ability to perform density estimation—modeling the probability distribution of potential outcomes based on input data. This capability is a key function in predictive analytics and estimation tasks within transportation systems. Specifically, “prediction” refers to forecasting future traffic conditions using historical and real-time data, while “estimation” focuses on determining the current status of the system, which often requires filling in data gaps or refining data accuracy. The ability to generate and evaluate multiple potential scenarios from DGMs is invaluable for effective traffic management. It allows transportation planners and engineers to anticipate and respond to a range of possible current and future conditions, thereby improving the efficiency and safety of the traffic network. By providing a probabilistic view of potential outcomes, DGMs empower decision-makers to accommodate the inherent variability and dynamic nature of traffic flows, significantly enhancing both the precision and reliability of transportation systems planning.
Building on the foundational understanding of DGMs and their application in traffic estimation and prediction, the following sections will discuss current research from three distinct scale levels: agent, link, and region. At the agent level, we will explore how these models are applied in individual vehicle and pedestrian prediction and estimation, critical for autonomous driving and safety systems. At the link level, the focus will shift to how DGMs are used for modeling traffic flows and speed on specific roadway segments, which is crucial for daily traffic management and dynamic routing. Finally, at the regional level, we will discuss the role of DGMs in shaping strategic planning and operational decisions by analyzing and predicting wide-area traffic patterns. Through this comprehensive examination, we aim to highlight the versatility and impact of DGMs in addressing the complexities of modern transportation challenges, thereby showcasing their essential role in the evolution of smart mobility solutions.
3.2.1   Agent-level Analysis
Agent-level analysis focuses on understanding and predicting the behaviors of individuals, such as vehicles, cyclists, or pedestrians. This detailed analysis is crucial for various applications like autonomous vehicle navigation, pedestrian safety systems, and dynamic routing. A significant portion of research in agent-level traffic analysis using DGMs primarily emphasizes learning the underlying distributions of traffic data (Roy et al., 2019; Zhao et al., 2020; Hegde et al., 2020; Zhou et al., 2021; Wang and He, 2021; Jagadish et al., 2021, 2022; Chen et al., 2022c; Hsu et al., 2023; Westny et al., 2024). This learning process is essential because it allows models to capture the intricate patterns and correlations within data that traditional models might miss. For example, Gómez-Huélamo et al. (2022) explored the use of attention mechanisms with GANs for vehicle trajectory prediction. The model generates feasible and realistic trajectories by considering both social interactions and the physical constraints of the road network. Similarly, Guo et al. (2023) developed the Map-Enhanced GAN (ME-GAN) for predicting autonomous vehicle trajectories. It leverages vehicle motion features, traffic state, and high-definition map constraints within the GAN framework to enhance prediction accuracy. Zhong et al. (2022) introduced the Spatial-Temporal Generative Model (STGM) to predict trajectories in autonomous driving environments. STGM combines a CVAE with dual spatial and temporal encoders to extract features from the driving environment. EquiDiff, proposed by Chen et al. (2023) is a conditional equivariant diffusion model that integrates a denoising diffusion probabilistic model with an SO(2)-equivariant transformer, effectively managing uncertainties in precise vehicle trajectory predictions.
Another critical aspect of using DGMs in agent-level analysis is the ability to perform density estimation of outputs. This capability allows for generating multiple potential outcomes, offering a comprehensive probabilistic view that supports robust decision-making under uncertainty. Research has increasingly focused on this feature, enabling traffic managers and autonomous vehicle systems to adapt dynamically to varied and unpredictable conditions (Feng et al., 2019; Zhao et al., 2019; Cheng et al., 2020; Neumeier et al., 2021; Wang et al., 2020a; Eiffert et al., 2020; Rossi et al., 2021b; Choi et al., 2021b; Li et al., 2021; Rossi et al., 2021a; Oh and Peng, 2022; Xing et al., 2022; Jagadish et al., 2022; De Miguel et al., 2022; Gui et al., 2022; Yao et al., 2023; Tang et al., 2024; Lin et al., 2024). For example, Gupta et al. (2018) proposed the Social GAN to predict socially acceptable future trajectories for pedestrians in crowded scenes. The GAN model captures complex dynamics and social behaviors, generating a wide range of potential trajectories. Similarly, Kosaraju et al. (2019) introduced Social-BiGAT, a trajectory forecasting model that integrates Bicycle-GAN with Graph Attention Networks (GATs) and LSTMs. This model generates multimodal predictions of pedestrian trajectories by leveraging complex social interactions and physical context cues. In another study, Sun et al. (2021a) introduced a normalizing flow-based prediction model designed to model the exact probability distribution of future human trajectories. Kim et al. (2021) enhanced the accuracy of ego vehicle trajectory prediction using a driving style-based CVAE model. The proposed model integrates a DeepConvLSTM network for driving style recognition from in-vehicle sensor data, improving trajectory prediction. Wang et al. (2020c) presented a GAN-based framework for vehicle trajectory prediction that includes a novel auto-encoder social convolution and a recurrent social mechanism to capture both spatial and temporal interactions among vehicles. Expanding into air traffic, Wu et al. (2022b) proposed a GAN-based long-term 4D trajectory prediction method that generates multiple trajectories considering the spatiotemporal dynamics of flight paths. In a multi-agent context, Vishnu et al. (2023) incorporated traffic states into Transformer, GAN, and CVAE models to enhance multi-agent trajectory predictions. These models generate plausible trajectory outcomes that are contextually relevant and highly predictive. Salzmann et al. (2020) developed Trajectron++, an advanced VAE-based model that forecasts the trajectories of multiple agents in dynamic environments. Trajectron++ integrates a CVAE with Gaussian Mixture Models (GMMs) within a graph-based recurrent neural network framework. This structure allows Trajectron++ to generate multiple potential trajectories in a probabilistic manner, providing a comprehensive representation of possible future scenarios. Li et al. (2019a) proposed the Conditional Generative Neural System (CGNS), which combines the CVAE with GAN to generate feasible, realistic, and diverse future trajectories for multiple agents by leveraging both static context and dynamic interactions. Bhattacharyya et al. (2019) introduced Conditional Flow VAE (CF-VAE), which uses conditional normalizing flows to improve the forecast of future states by capturing complex, multi-modal conditional distributions. Li et al. (2023c) developed a framework combining CVAE and conditional diffusion models for multi-modal vehicle trajectory prediction. This approach addresses the challenge of predicting highly uncertain future vehicle trajectories in urban environments by first generating trajectories with CVAE and refining them using a diffusion model. Similarly, Liu et al. (2024) aimed to predict the distribution of endpoints of multi-agent trajectories using denoising diffusion and Transformer models, capturing both the spatio-temporal dynamics and the intrinsic intent of vehicles to enhance overall prediction quality. These examples illustrate the diverse and robust capabilities of DGMs in agent-level analysis, enabling precise and reliable predictions that are critical for the development and operation of intelligent transportation systems.
3.2.2   Link-level Analysis
Effective traffic management at the link level is essential for maintaining smooth traffic flow and ensuring safety on individual road links within transportation networks. This process includes accurate predictions and estimations of traffic conditions such as flow, speed, and travel time. DGMs play a pivotal role at this scale by providing advanced insights that support real-time traffic management, routing optimization, and congestion control. By leveraging the power of DGMs, traffic managers can respond more effectively to changing conditions on each link and anticipate future variations with greater precision. As discussed earlier, one of the benefits of using DGMs in estimation and prediction tasks is their ability to learn the characteristics and distributions from training data, enabling them to generate realistic and reliable data for effective traffic decision-making (Lin et al., 2018; Chen et al., 2018; Zhang et al., 2019a, b; Impedovo et al., 2019; Zang et al., 2019; Xu et al., 2020a; Li et al., 2020b; Yu et al., 2020a; Aibin et al., 2021; Sun et al., 2021b; Zhang et al., 2021a; Zhou et al., 2020; Song et al., 2021; Wang et al., 2022a; Zhao et al., 2022; Mo et al., 2022a; Xu et al., 2022; Khaled et al., 2022). For example, Liang et al. (2018) introduced a GAN-based model that incorporates two LSTM networks and traffic-flow theory for spatial-temporal traffic state estimation. The GAN generates realistic traffic-state values by learning from the underlying data distributions. Similarly, Li et al. (2019b) presented DeepGTT, a VAE-based model designed to predict travel time distributions by integrating real-time traffic data and spatial features. The VAE enables the model to generate realistic travel time distributions even under conditions of data sparsity and variability. In another study, Xu et al. (2020b) developed the Graph Embedding GAN (GE-GAN), which selects adjacent links to estimate road traffic speed and volume more accurately. To be more specific, it uses Graph Embedding (GE) techniques to represent the spatial characteristics of road networks, while a Wasserstein GAN (WGAN) is employed to generate traffic state data. Jin et al. (2022) proposed the PL-WGAN, integrating the advantages of multiple deep learning units, including WGAN, GCN, RNN, and attention mechanisms, for predicting short-term traffic speed in urban networks. The WGAN in this framework generates realistic traffic speed data by learning from historical patterns and control parameters. Yu et al. (2020b) introduced a framework combining clustering and generative models to predict taxi hotspots. This model, which integrates LSTM with CGAN, learns the spatiotemporal distribution of taxi hotspots to generate accurate predictions. Moreover, Zhou et al. (2020) proposed a Bayesian framework that combines VAEs with GNNs for robust traffic prediction. The model benefits from the generative capabilities of VAEs and Normalizing Flow to capture multimodal traffic data distributions, addressing the uncertainty and complexity of road sensor networks. Recently, Mo et al. (2022b) successfully combined normalizing flows, GANs, and physics-informed deep learning to generate realistic traffic data while quantifying uncertainty. The model uses the normalizing flow model as the generator within the GAN framework to explicitly estimate data likelihood. Additionally, physics-informed deep learning is employed to integrate traffic domain knowledge, ensuring that the model adheres to the laws of traffic flow in data estimation.
While many studies using DGMs focus on learning complex data distributions to generate accurate and realistic outcomes, another critical aspect is the density estimation of outputs. This feature is less explored but crucial for providing a probabilistic view of potential outcomes, enhancing decision-making under uncertainty (Koochali et al., 2019; Arnelid et al., 2019; Rasul et al., 2020, 2021; Wen et al., 2023). For instance, Koochali et al. (2019) presented ForGAN, a probabilistic forecasting model that uses conditional GANs to predict future sensory data by learning the full conditional probability distribution from historical data. This approach employs KL Divergence to measure and refine the accuracy of predicted distributions, ensuring they closely match true data distributions. Rasul et al. (2020) developed an autoregressive deep learning model for multivariate probabilistic time series forecasting using conditioned normalizing flows. The model captures complex dependencies between multiple time series by modeling the joint distribution of future observations conditioned on past data, significantly improving prediction accuracy and uncertainty estimation. Rasul et al. (2021) introduced TimeGrad, which applies a denoising diffusion model to the same problem. TimeGrad samples from the data distribution at each time step by estimating its gradient and transforming noise into meaningful data through a Markov chain using Langevin sampling. Similarly, Wen et al. (2023) developed DiffSTG, a framework for probabilistic Spatio-Temporal Graph (STG) forecasting that combines Spatio-Temporal GNNs (ST-GNNs) with DDPMs. DiffSTG captures both spatial and temporal dependencies in STG data while modeling inherent uncertainties. This model addresses the inefficiencies of TimeGrad in long-term forecasting by using a non-autoregressive approach to predict multiple future horizons simultaneously. In another study, Feng et al. (2024) introduced the Latent Diffusion Transformer (LDT) for high-dimensional multivariate probabilistic time series forecasting. LDT employs a latent space approach combined with a diffusion-based conditional generator and a symmetric statistics-aware autoencoder to enhance the expressiveness and manageability of forecasting complex time series data. These studies demonstrate the robust capabilities of DGMs in link-level analysis, offering essential tools for predicting and managing traffic conditions on road networks, which further contributes to more efficient and responsive traffic management systems.
3.2.3   Region-level Analysis
At the regional level, traffic management includes estimating and predicting traffic conditions across large geographic areas encompassing multiple road networks and transportation systems. This type of analysis is essential for understanding broader traffic patterns and making informed decisions that impact urban planning, resource allocation, and emergency response strategies. Key tasks at the region level include predicting and estimating average speeds across the network, estimating origin-destination (OD) flows to understand travel patterns, forecasting traffic demand to anticipate future needs, analyzing overall traffic flow to identify congestion areas, and performing crash analysis to enhance network safety. The complexity of managing traffic at the regional level arises from the need to integrate data from diverse sources and understand the dynamic interactions between various components of the transportation network. From this perspective, DGMs are particularly well-suited due to their capability to learn from extensive datasets, capture complex spatial and temporal patterns, and model uncertainties of output results. This enables DGMs to provide accurate and robust estimations and predictions crucial for strategic decision-making and enhancing the efficiency and safety of transportation systems. Many studies have applied DGMs in the region-level traffic estimation and prediction (Saxena and Cao, 2019; Zhang et al., 2020b; Li et al., 2020a; Kakkavas et al., 2021; Wu et al., 2021; Feng et al., 2021; Kakkavas et al., 2021; Naji et al., 2021; Wang et al., 2022b; Li et al., 2022c, b; Lin et al., 2023b; Zhang et al., 2023a; Zarei et al., 2024; Li et al., 2024b; Shao et al., 2024). For example, Yu and Gu (2019) proposed a deep neural network architecture called the Graph Convolutional Generative Autoencoder (GCGA) to address real-time speed estimation problems. In GCGA, the Graph Convolutional Network (GCN) extracts spatial characteristics, and an autoencoder-based GAN uses these features to generate traffic speed maps from incomplete data. Yu et al. (2019) developed a deep generative model framework that integrates a modified DBSCAN algorithm and a CGAN model to predict taxi-passenger demand. The model accounts for spatial, temporal, and external dependencies of input data, providing accurate demand predictions. In another study, Wang et al. (2020b) introduced the SeqST-GAN, a sequence-to-sequence GAN model comprising LSTM and CNN for multi-step urban crowd flow prediction. This model enhances long-term prediction accuracy by treating crowd flow data as "image frames" and incorporating contextual features. Huang et al. (2022b) presented the Dynamic Multi-Graph Convolutional Network with GAN framework (DMGC-GAN) for predicting OD-based ride-hailing demand. DMGC-GAN effectively addresses high sparsity and complex spatio-temporal correlations in OD demand data by leveraging dynamic, directed OD graphs within a GAN framework. Furthermore, Lin et al. (2023b) developed Curb-GAN, a conditional GAN-based model for estimating urban traffic conditions under various travel demand scenarios. Curb-GAN integrates dynamic convolutional layers and self-attention mechanisms to capture spatial and temporal correlations, producing accurate and realistic traffic estimations. Yuan et al. (2023) introduced a model based on the DDPM for traffic matrix estimation. The DDPM learns the distribution of traffic matrices and improves estimation accuracy by incorporating a preprocessing module and framing the estimation task as a gradient-descent optimization problem. These studies illustrate how DGMs can effectively manage and predict traffic conditions at the regional level, providing valuable insights for strategic planning and operational decision-making across large transportation networks.
3.3 DGM for Unsupervised Representation Learning
Unsupervised representation learning involves training a model to learn useful features or representations of the data without requiring labeled inputs. This method is valuable in environments where labeled data is sparse or expensive, which is often the case in transportation datasets. DGMs, such as VAEs are adept at discovering intricate structures in unlabeled data by learning to compress data and reconstruct it back into the original space. In other words, DGMs learn to encode data into a compact, latent space—a lower-dimensional representation of the original data—which preserves much of the information but in a more compressed form. The latent vector is a compact representation of the input data that captures its most critical features. These vectors are the output of the model’s encoder component and serve as a compressed knowledge base of the data, containing the essential information needed to reconstruct or generate new data points. In transportation research, the manipulation and analysis of these latent vectors allow researchers to infer traffic patterns, and anomalies that might not be apparent in the high-dimensional original data.
One primary research direction is feature extraction and classification. It gains popularity in the field of transportation when employing DGMs, particularly because of their potent capability to discern and categorize complex, multi-dimensional data without direct supervision. Many studies have used the latent space in various classification tasks such as transportation mode identification, driving behavior analysis, and anomaly detection (Krajewski et al., 2018; Rákos et al., 2021, 2020; Yao and Bekhor, 2022; Rakos et al., 2021; De Miguel et al., 2022; Xie et al., 2023; Santhosh et al., 2021; Islam et al., 2021; Krajewski et al., 2019; Ding et al., 2019; Yuan et al., 2023; Kim et al., 2021). For example, Gupta et al. (2018) proposed Social GAN and explored how variations in the latent variables of a generative model influence the generated human trajectories. They discovered that specific directions in the latent space correspond to changes in trajectory direction and speed, providing insights into controlling the types of behaviors the model can generate. Zhang et al. (2022c) introduced the Dirichlet VAE (DirVAE) to identify transportation modes from GPS trajectory data. The latent space of the DirVAE visualizes and classifies different transportation modes, demonstrating the utility of latent space in distinguishing between various modes of transportation. In another study, Boquet et al. (2020) used the VAE to represent complex, high-dimensional traffic data. The VAE learns a low-dimensional latent space representation that captures the underlying patterns and dependencies within the traffic data. This latent space helps visualize the distribution of different traffic data types, such as normal and anomalous patterns. Neumeier et al. (2021) proposed an unsupervised model based on the VAE architecture to predict vehicle trajectories with an interpretable latent space. The latent space can be analyzed to understand and predict lane-changing maneuvers, enhancing the model’s utility in studying vehicle behaviors. Chen et al. (2021c) developed a method for generating realistic traffic flow data using GANs with semantic latent code manipulation. By exploring and manipulating the semantic representations in the latent space, their model generates traffic flow data that accurately reflects realistic patterns and conditions. Moreover, Santhosh et al. (2021) proposed a novel approach combining CNN and VAE to classify vehicle trajectories and detect anomalies. This hybrid model addresses the challenges of classifying time-series data of varying lengths and identifying anomalies such as lane violations, sudden speed changes, and vehicles moving in the wrong direction. The visualization of the latent space in their model provides a clear understanding of how the VAE encodes trajectory data and distinguishes between normal and anomalous patterns. These studies highlight the effectiveness of DGMs in feature extraction and classification tasks within transportation. By fully utilizing the latent space, researchers can find intricate patterns and dependencies, facilitating more accurate and insightful analysis of transportation data.
3.4 Summary
This section reviewed the significant role of DGMs in transportation research, focusing on their applications in data generation, estimation and prediction, and unsupervised representation learning. In data generation, DGMs are invaluable for producing synthetic datasets that replicate real-world traffic conditions. This capability is essential for training models, testing new scenarios, and supplementing real data, particularly when data collection is challenging or costly. DGMs can generate diverse traffic scenarios, realistic trajectories, and even complete missing data, thereby ensuring comprehensive and robust datasets for traffic analysis and management. For estimation and prediction, DGMs excel in modeling the uncertainties and variations inherent in traffic data. By learning from extensive datasets, they provide probabilistic forecasts of traffic conditions and enable effective decision-making for traffic management and planning. These models support the anticipation and adaptation to probable future traffic patterns, contributing to more efficient and responsive transportation systems. Lastly, in unsupervised representation learning, DGMs facilitate the analysis of high-dimensional traffic data by encoding it into compact, interpretable latent spaces. This enables the classification of transportation modes, analysis of driving behaviors, and detection of anomalies, providing deeper insights without the need for labeled data. The ability of DGMs to reveal hidden structures within data is particularly valuable for understanding and improving transportation systems.
Despite these advances in transportation studies, future research must address several key challenges. Standardizing evaluation methods is crucial to ensure consistency and comparability across studies. Future studies should also consider the dynamic characteristics of traffic data, which vary significantly across different spatial and temporal contexts. Additionally, understanding and incorporating the causal relationships within traffic data, as well as addressing ethical and privacy concerns associated with synthetic data generation, are critical areas that require further exploration. These challenges, along with other emerging trends and opportunities in the use of DGMs for transportation research, will be discussed in more detail in the following Section 5.
Tutorial
In this section, we present practical examples of how DGMs can be applied in transportation research. To reach a broader audience, we provide two hands-on tutorials: 1) Generating Household Travel Survey Data in Section 4.1, and 2) Generating Highway Traffic Speed Contour (Time-Space Diagram) in Section 4.2. Importantly, all data and code used in these tutorials—including preprocessing scripts, model training, and inference code, and pre-trained model parameters—are available in our GitHub repository: https://github.com/benchoi93/DGMinTransportation
This tutorial aims to equip researchers with the necessary tools to explore the application of DGMs in transportation studies, rather than identifying the top-performing models within each research domain. To this end, we do not perform numerical or analytical evaluations of model performance. Instead, the focus is on providing qualitative insights, which is a common practice in DGM research. For readers interested in rigorous performance evaluations, including numerical or analytical assessments, we suggest referring to the existing literature. Nevertheless, the code provided in our repository offers a robust foundation for initiating DGM-based research, allowing researchers to extend these models for specific transportation study requirements.
4.1 Generating Household Travel Survey Data
Household Travel Survey (HTS) data provides a comprehensive overview of individuals’ daily travel patterns, modes of transportation, and trip purposes within sampled households. These datasets often include sociodemographic and household-level characteristics, making them indispensable for transportation planners, policymakers, and researchers. HTS data is crucial for understanding mobility patterns, forecasting travel demands, and informing infrastructure developments, public transit planning, and sustainable transportation initiatives (Zhang and Li, 2022). However, traditional methods of collecting HTS data, such as face-to-face interviews or telephone surveys, typically yield sample sizes representing only 1–5% of the population in a given region (Stopher and Greaves, 2007), raising concerns about their representativeness and accuracy. Additionally, low sample rates, high survey costs, and safety concerns of interviewers further complicate the data collection process. While some countries, like Singapore, have adopted the online HTS collection system, questions about participant fidelity remain. These limitations underscore the need for innovative approaches, such as synthetic HTS data generation, to more accurately represent population-wide travel behaviors.
Advanced artificial intelligence techniques, particularly DGMs, have emerged as a promising solution to address the constraints of traditional HTS data collection. While prior studies have explored the use of GAN or VAE for synthetic data generation (Kim and Bansal, 2023; Garrido et al., 2020; Borysov et al., 2019), other DGMs remain underexplored. Therefore, this section offers a tutorial on generating a simplified version of HTS data using DGMs, demonstrating their potential to overcome traditional data collection limitations and laying the groundwork for future research and applications in this field.
4.1.1   Data and Preprocessing
In this tutorial, we applied DGMs to the 2010 Household Travel Survey data from South Korea. This dataset offers comprehensive details on household travel patterns, including trip purposes, transportation modes, and travel times across various cities and districts in the region. It serves as a valuable resource for investigating urban mobility dynamics. While the original HTS dataset covers the entire country, this study only focuses specifically on the Seoul Metropolitan Area, as illustrated in Figure 7. The Seoul Metropolitan Area encompasses Seoul, Incheon, and Gyeonggi-do located in the northwest region of South Korea. With a population of approximately 26 million, this area represents more than half of the country’s total population.
Figure 7: Seoul Metropolitan Area
The original household travel survey includes sociodemographic data and completed travel data for the households. For simplicity, we exclude the sociodemographic data and focus solely on travel data. An example of the HTS travel data used in this study is shown in Figure 8. This travel data that are composed of three different trips can be transformed into the tabular data as the Table 2. To simplify the problem, we selected the types of origin, activity, mode choice, and destination as target features for data generation. Table 3 provides a detailed description of the chosen features. In our tutorial example, each DGM is designed to generate each row that contains the data in Table 3 by learning the joint distribution of the selected features. The data includes origin and destination types divided into five categories: home, work, school, transfer points, and other locations. Transportation modes are categorized into nine discrete integers: 0 for the start of the day, 1 for the end of the day, 2 for staying at a location, and 3 through 8 representing various modes of transport such as walking, public transit, driving (driver and passenger), cycling, and taxis. Lastly, activity types include eight categories: work, education, leisure, shopping, and escort activities. It is important to note that the goal of this tutorial is not generating the complete daily travel, but rather focusing on individual trips. This allows the model to capture the relationships between features. During model evaluation, we present qualitative assessments with marginal distributions of each feature. This qualitative approach prioritizes interpretability and understanding over direct numerical performance metrics. Nevertheless, the model is trained to capture the joint probability distribution, emphasizing its ability to generate realistic combinations of attributes.
Figure 8: Example daily travel in the HTS data
Table 2: Trip chain data generated from example daily travel in Figure 8
Origin Type Origin District Num. Departure Time Activity Type Mode Type Destination Type Destination District Num. Arrival Time
1 Home Gyeonggi-1 00:00 AM Start Stay Home Gyeonggi-1 08:30 AM
2 Home Gyeonggi-1 08:30 AM Travel Transit Work Place Seoul-5 09:00 AM
3 Work Place Seoul-5 09:00 AM Work Stay Work Place Seoul-5 06:00 PM
4 Work Place Seoul-5 06:00 PM Travel Taxi Others Gyeonggi-2 06:30 PM
5 Others Gyeonggi-2 06:30 PM Shopping Stay Others Gyeonggi-2 08:15 PM
6 Others Gyeonggi-2 08:15 PM Travel Walk Home Gyeonggi-1 08:30 PM
7 Home Gyeonggi-1 08:30 PM End Stay Home Gyeonggi-1 12:00 PM
Table 3: Details of each attribute in generated data
Features Range of Value Description on Each Value
Origin Type Discrete Integer (0$\sim$5) 0: Home, 1: Work Place, 2: School 3: Transfering 4: Other
Activity Type Discrete Integer (0$\sim$8) 0: Home, 1: Travel, 2: Work, 3: Education, 4: Academy, 5: Leisure, 6: Shopping, 7: Escort, 8: Other
Mode Type Discrete Integer (0$\sim$8) 0: Start, 1: End, 2: Stay, 3: Walk, 4: Transit, 5: Car (driver), 6: Car (passenger), 7: Bicycle, 8: Taxi
Destination Type Discrete Integer (0$\sim$5) 0: Home, 1: Work Place, 2: School 3: Transfering 4: Other
Most DGMs are typically designed to operate in continuous spaces, learning to map the probability distribution of data. This design presents challenges when applied to discrete datasets, such as HTS data. When dealing with discrete variables, the probability distribution only has values at specific discrete points, while the probability is zero everywhere else. Additionally, to normalize the probabilities to sum to 1, the probability at these specific data points can become infinitely large. This causes DGMs to assign infinite likelihood to certain points, which makes the model unstable. To address this issue, adding noise to the discrete variables, i.e., relaxed one hot encoding, was suggested from the literature. (Jang et al., 2016) presented Categorical VAEs, which incorporate noise from Gumbel-Softmax distribution. (Maddison et al., 2016) showed a similar approach using Concrete distribution to approximate the discrete value into continuous value.
In this tutorial, we use a simple dequantization technique to address the challenges of generating discrete data using DGMs. Rather than relying on complex methods such as Gumbel-Softmax to model discrete or categorical distribution, we add uniform random noise to each discrete data sample before training the model, effectively transforming the discrete data into continuous space. After data generation using DGMs, we apply a floor operation to convert it back to the discrete form. Mathematically, this process transforms the generation of a discrete value $\mathbf{x}_{i}$ into generating a continuous value within the range $[\mathbf{x}_{i},\mathbf{x}_{i}+1)$. Consequently, computing the probability $p(\mathbf{x})$ is expressed as:
[…]
where $q(\mathbf{u}|\mathbf{x})$ is the noise distribution. Since we assume it to be uniform, this can be rewritten as:
$p(\mathbf{x})=\mathbb{E}_{\mathbf{u}\sim U\left[0,1\right)^{D}}\left[p(\mathbf{x}+\mathbf{u})\right],$ (4.2)
where $U\left[0,1\right)^{D}$ is a $D$-dimensional (data dimension) random uniform vector. This method effectively smooths the distribution and ensures stable training and generation of discrete variables.
This approach works well across various DGMs, as we demonstrate in this tutorial. The Python implementation of this dequantization process is included in the accompanying codebase. All five DGMs presented in this tutorial follow the same dequantization process, ensuring consistency and comparability across models.
1 def dequantize_data(data):
2
3 data_continuous = data.copy().astype(float)
4 size = data_continuous.shape[0]
5
6 for column in data_continuous.columns:
7 data_continuous[column] += np.random.uniform(0, 1, size=size)
8 return data_continuous
4.1.2   Variational Autoencoder
The VAE loss function consists of two primary components as shown in Equation (2.11): the reconstruction loss and the regularization loss. The reconstruction loss measures how well the model can reproduce the original data. Since the encoder and decoder are both deterministic functions when input is given, the reconstruction loss term, $\mathbb{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]$, can be re-written as $\mathbb{E}_{\mathbf{x}}\left[\log p_{\theta}(\mathbf{\hat{x}}|\mathbf{x})\right]$. If we assume that this distribution has a Gaussian form:
$\log p_{\theta}(\mathbf{\hat{\mathbf{x}}}|\mathbf{\mathbf{x}})\propto\log e^{-|\mathbf{x}-\hat{\mathbf{x}}|^{2}}=-|\mathbf{x}-\hat{\mathbf{x}}|^{2},$ (4.3)
then maximizing $\mathbb{E}_{\mathbf{x}}\left[\log p_{\theta}(\mathbf{\hat{x}}|\mathbf{x})\right]$ corresponds to minimizing the square loss between $\mathbf{x}$ and $\mathbf{\hat{x}}$. The regularization loss, on the other hand, ensures that the induced latent space from the encoder (i.e., the posterior distribution) follows the same distribution as the prior distribution. However, in practical implementations, the regularization term often requires simplifying assumptions. In this case, we assume that both the posterior distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ and the prior distribution $p(\mathbf{z})$ are Gaussian distribution. Under these assumptions, the Kullback-Leibler (KL) divergence between the two Gaussian distributions, considering D-dimensional vector $\mathbf{z}$ can be simplified into a more computationally manageable form as:
$\begin{split}D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)&=D_{KL}\left(\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})||\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})\right)\\ &=\frac{1}{2}\left(\mathrm{Tr}(\boldsymbol{\Sigma}_{2}^{-1}\boldsymbol{\Sigma}_{1})+(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}_{2}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})-D+\log\frac{\det\boldsymbol{\Sigma}_{2}}{\det\boldsymbol{\Sigma}_{1}}\right).\\ \end{split}$ (4.4)
where $(\boldsymbol{\mu}_{1},\mathbf{\Sigma}_{1})$, and $(\boldsymbol{\mu}_{2},\mathbf{\Sigma}_{2})$ are the mean vectors and covariance matrices for $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $p(\mathbf{z})$, respectively. We can further simplify the equations by setting the prior distribution, $p(\mathbf{x})$, to follow standard Gaussian distribution, i.e., $\boldsymbol{\mu}_{2}=0$, $\boldsymbol{\Sigma}_{2}=\mathbf{I}$, and the posterior distribution, $q_{\phi}(\mathbf{z}|\mathbf{x})$, follows zero-mean Gaussian distribution with diagonal covariance matrix, i.e., $\boldsymbol{\mu}_{1}=0$, and $\boldsymbol{\Sigma}_{1}=\boldsymbol{\sigma}_{1}^{2}\mathbf{I}$:
$\begin{split}D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)&=\frac{1}{2}\left(\mathrm{Tr}(\boldsymbol{\Sigma}_{1})+\boldsymbol{\mu}_{1}^{\top}\boldsymbol{\mu}_{1}-D-\log\det\boldsymbol{\Sigma}_{1}\right)\\ &=\frac{1}{2}\sum_{i=1}^{D}\left(\sigma_{1,i}^{2}+\mu_{1,i}^{2}-1-\log\sigma_{1,i}^{2}\right),\\ \end{split}$ (4.5)
where $\mu_{1,i}$ is $i$-th element in $\boldsymbol{\mu}_{1}$ and $\sigma_{1,i}$ is $i$-th element in $\boldsymbol{\sigma}_{1}$. The corresponding code implementation of the VAE loss is stated below. We use $\log\boldsymbol{\sigma}_{1}$ instead of $\boldsymbol{\sigma}_{1}$ for numerical stability, and the variable logvar corresponds to $2\log\boldsymbol{\sigma}_{1}$. The variable mu corresponds to $\boldsymbol{\mu}_{1}$. Figure 9 illustrates the structure of the neural network for this section. Figure 9 (a) shows the structure of the encoder which encodes the inputs to $D$-dimensional vectors for both Mean and Variance and we used $D=64$, and Figure 9 (b) shows the structure of the decoder which generates the sample from the latent vector. The code implementation is as follows:
Input (x) Linear (4 $\to$ 512) ReLU Linear (512 $\to$ 512) ReLU Linear (512 $\to$ 64) (mu_layer) Linear (512 $\to$ 64) (log_var_layer) Mean (mu) Variance (logvar)
(a) Network structure of encoder
Latent Vector (z) Linear (64 $\to$ 512) ReLU Linear (512 $\to$ 512) ReLU Linear (512 $\to$ 4) Sample (x_reconstructed)
(b) Network structure of decoder
Figure 9: Neural network structures of VAE in HTS data
1 class VAE(nn.Module):
2 def __init__(self, input_dim, hidden_dim, z_dim):
3 super(VAE, self).__init__()
4 # Encoder
5 self.encoder = nn.Sequential(
6 nn.Linear(input_dim, hidden_dim),
7 nn.ReLU(),
8 nn.Linear(hidden_dim, hidden_dim),
9 nn.ReLU()
10 )
11 self.mu_layer = nn.Linear(hidden_dim, z_dim)
12 self.log_var_layer = nn.Linear(hidden_dim, z_dim)
13
14 # Decoder
15 self.decoder = nn.Sequential(
16 nn.Linear(z_dim, hidden_dim),
17 nn.ReLU(),
18 nn.Linear(hidden_dim, hidden_dim),
19 nn.ReLU(),
20 nn.Linear(hidden_dim, input_dim)
21 )
22
23 def encode(self, x):
24 h = self.encoder(x)
25 mu = self.mu_layer(h)
26 log_var = self.log_var_layer(h)
27 return mu, log_var
28
29 def reparameterize(self, mu, log_var):
30 std = torch.exp(0.5 * log_var)
31 eps = torch.randn_like(std)
32 return mu + eps * std
33
34 def decode(self, z):
35 return self.decoder(z)
36
37 def forward(self, x):
38 mu, log_var = self.encode(x)
39 z = self.reparameterize(mu, log_var)
40 x_reconstructed = self.decode(z)
41 return x_reconstructed, mu, log_var
42
43 [...]
44 vae_model = VAE(input_dim=input_dim, hidden_dim=hidden_dim, z_dim=z_dim)
45
46 [...] # in training loop
47 x_reconstructed, mu, log_var = vae_model(x) # calls forward function in VAE class
48 recon_loss = F.mse_loss(x_reconstructed, x, reduction=’sum’) # Equation (4.3)
49 kl_loss = -0.5 * torch.sum(1 + logvar - mu.pow(2) - logvar.exp()) # Equation (4.4)
50 loss = recon_loss + kl_loss
4.1.3   Generative Adversarial Networks
Unlike VAEs, GANs do not require complex derivations of the loss function. The code implementation of the Equations (2.13) and (2.14) can be written intuitively. In practice, this involves using binary cross-entropy loss. The discriminator tries to identify if the given input is from the real data set or the generated data set. On the contrary, the generator tries to overcome the identification from the discriminator. Similar to the VAE, we constructed the 3-layer neural network for (a) generator and (b) discriminator as shown in Figure 10. The noise vector (latent vector) is $D$-dimensional vector where $D=64$. The code implementation is as follows:
1 class Generator(nn.Module):
2 def __init__(self, input_dim, output_dim, hidden_dim):
3 super(Generator, self).__init__()
4 self.net = nn.Sequential(
5 nn.Linear(input_dim, hidden_dim),
6 nn.ReLU(),
7 nn.Linear(hidden_dim, hidden_dim),
8 nn.ReLU(),
9 nn.Linear(hidden_dim, output_dim)
10 )
11
12 def forward(self, z):
13 return self.net(z)
14
15 class Discriminator(nn.Module):
16 def __init__(self, input_dim, hidden_dim):
17 super(Discriminator, self).__init__()
18 self.net = nn.Sequential(
19 nn.Linear(input_dim, hidden_dim),
20 nn.ReLU(),
21 nn.Linear(hidden_dim, hidden_dim),
22 nn.ReLU(),
23 nn.Linear(hidden_dim, 1),
24 nn.Sigmoid()
25 )
26
27 def forward(self, x):
28 return self.net(x)
29
30 [...]
31 generator = Generator(input_dim=z_dim, output_dim=input_dim, hidden_dim=hidden_dim)
32 discriminator = Discriminator(input_dim=input_dim, hidden_dim=hidden_dim)
33
34 criterion = nn.BCELoss()
35 real_label = torch.ones(batch_size)
36 fake_label = torch.zeros(batch_size)
37
38 [...] # in training loop
39 z = torch.randn(batch_size, z_dim) # random number generator from Standard Gaussian Distribution (N(0,1))
40 fake_data = generator(z) # Figure 10 (b)
41
42 d_real = discriminator(real_data) # Figure 10 (a)
43 d_fake = discriminator(fake_data) # Figure 10 (a)
44 d_loss_real = criterion(d_real, real_label)
45 d_loss_fake = criterion(d_fake, fake_label)
46 [...] # update discriminator
47
48 z = torch.randn(batch_size, z_dim) # random number generator from Standard Gaussian Distribution (N(0,1))
49 fake_data = generator(z) # Figure 10 (b)
50 g_fake = discriminator(fake_data) # Figure 10 (a)
51 g_loss = criterion(discriminator(fake_data), real_label)
52 [...] # update generator
Inputs (real_data fake_data) Linear (4 $\to$ 512) ReLU Linear (512 $\to$ 512) ReLU Linear (512 $\to$ 4) Sigmoid Decision (d_real, d_fake, g_fake)
(a) Network structure of discriminator
Latent Vector (z) Linear (64 $\to$ 512) ReLU Linear (512 $\to$ 512) ReLU Linear (512 $\to$ 4) Sample (fake_data)
(b) Network structure of generator
Figure 10: Neural network structures of GAN in HTS data
4.1.4   Normalizing Flows (Flow-based Generative Models)
We use a flow-based generative model inspired by RealNVP (Dinh et al., 2016), which leverages affine coupling layers and alternating masking strategies to transform complex data distributions into a standard normal distribution. For simplicity, and since the data dimensionality is not too large, the model in this tutorial does not include permutation layers between coupling layers and batch normalization that were proposed in the original RealNVP paper. A more detailed implementation of RealNVP will be discussed in Section 4.2.4. Despite these simplifications, our model maintains the core idea of using invertible affine transformations to compute the log-likelihood and optimize the data distribution. The training objective is to maximize the likelihood of the observed data under this transformation, which can be expressed as minimizing the negative log-likelihood. The original loss function for the normalizing flow is the same as stated in Equation (2.21). For the affine coupling layers, we use the same equations as described from Equation (2.23) to Equation (2.26). We use much simpler networks for scaling and translation network as shown in Figure 11. The code implementation is as follows:
Inputs (x_masked) Linear (4 $\to$ 128) ReLU Linear (128 $\to$ 128) ReLU Linear (128 $\to$ 4) Tanh Scale (s)
(a) Network structure of scale net
Inputs (x_masked) Linear (4 $\to$ 128) ReLU Linear (128 $\to$ 128) ReLU Linear (128 $\to$ 4) Translation (t)
(b) Network structure of translation net
Figure 11: Neural network structures of normalizing flow in HTS data
1 class AffineCoupling(nn.Module):
2 [...]
3 # provide calculation of log_det_jacobian using scale and translation functions
4 def forward(self, x):
5 x_masked = x * self.mask
6 s = self.scale_net(x_masked) * (1 - self.mask)
7 t = self.translate_net(x_masked) * (1 - self.mask)
8 y = x_masked + (1 - self.mask) * (x * torch.exp(s) + t) # Equation (2.23)
9 log_det_jacobian = torch.sum(s, dim=1) # Equation (2.26)
10 return y, log_det_jacobian
11 [...]
12
13 class NormalizingFlow(nn.Module):
14 def __init__(self, input_dim, hidden_dim, num_layers):
15 super(NormalizingFlow, self).__init__()
16 self.layers = nn.ModuleList()
17 mask = self.create_mask(input_dim, even=True)
18 for i in range(num_layers):
19 self.layers.append(AffineCoupling(input_dim, hidden_dim, mask))
20 # Alternate the mask for the next layer
21 mask = 1 - mask
22
23 def forward(self, x):
24 """
25 Forward pass through the flow: transform x to z and compute log_prob.
26 """
27 log_det_jacobian = torch.zeros(x.size(0))
28 for layer in self.layers:
29 x, ldj = layer(x)
30 log_det_jacobian += ldj
31 # Compute log_prob under base distribution
32 log_prob_z = self.base_log_prob(x)
33 log_prob = log_prob_z + log_det_jacobian # Equation (2.20)
34 return log_prob
35 [...]
36 log_prob = flow_model(x)
37 loss = -torch.mean(log_prob) # Equation (2.21)
4.1.5   Diffusion Models
In the tutorial code, we used the Denoising Diffusion Probabilistic Model (DDPM) as the reference for generating data (Ho et al., 2020). Instead of using the traditional U-Net architecture for noise prediction, we implemented a simpler neural network structure, as shown in Figure 12. The U-Net architecture is commonly used in diffusion models due to its powerful ability to capture multi-scale features in image data. However, in this tutorial, we implemented a simpler neural network to reduce computational complexity and facilitate a focus on the fundamental concepts of diffusion models, such as stepwise noise estimation and progressive denoising. The loss function for training is calculated based on Equation (2.39) presented above, which measures the model’s ability to predict noise accurately at each step of the diffusion process. The code implementation is as follows:
Inputs (x_t) Inputs (t) Linear (4+128 $\to$ 256) ReLU Linear (256 $\to$ 256) ReLU Linear (256 $\to$ 4) Noise (noise_pred)
Figure 12: Neural network structure of DDPM in HTS data
1 class DiffusionModel(nn.Module):
2 [...]
3
4 def forward_diffusion_sample(x0, t):
5 noise = torch.randn_like(x0)
6 sqrt_alpha_prod = sqrt_alphas_cumprod[t].view(-1, 1)
7 sqrt_one_minus_alpha_prod = sqrt_one_minus_alphas_cumprod[t].view(-1, 1)
8 x_t = sqrt_alpha_prod * x0 + sqrt_one_minus_alpha_prod * noise
9 return x_t, noise
10 [...]
11 t = torch.randint(0, T, (batch_size,), device=device).long()
12 x_t, noise = forward_diffusion_sample(x0, t)
13 noise_pred = diffusion_model(x_t, t)
14 loss = nn.functional.mse_loss(noise_pred, noise) # Equation (2.39)
4.1.6   Score-based Generative Model
The original objective function of the Score-based Model is defined as Equation (2.28). Here we show how the transformation to Equation (2.29) proceeds when we are using the normal distribution for the noise $q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)$. In this case, the equation can be simplified as shown follows:
$\begin{split}&\left|\left|{\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)-s_{\theta}(\tilde{\mathbf{x}},\sigma)}\right|\right|^{2}_{2}=\left|\left|{\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)}{\sigma^{2}}+s_{\theta}(\tilde{\mathbf{x}},\sigma)}\right|\right|^{2}_{2},\\ \end{split}$ (4.6)
since,
$\begin{split}&q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)^{2}}{2\sigma^{2}}\right),\\ &\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)^{2}}{2\sigma^{2}}-\log\left(\sqrt{2\pi\sigma^{2}}\right),\\ &\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)}{\sigma^{2}}.\end{split}$ (4.7)
where the term $\tilde{\mathbf{x}}-\mathbf{x}$ corresponds to the variable noise. We use a neural network with three linear layers and ReLU activations to obtain the estimated score as shown in Figure 13. In this tutorial, we train the NCSN using sequential sigma noise levels rather than randomly mixing them. This approach diverges from the typical method of random sampling of sigma values but was chosen to provide a clearer illustration of the overall training process. Given the relatively low dimension of the dataset, the sequential method is also feasible. The code implementation is as follows:
1 class ScoreNetwork(nn.Module):
2 def __init__(self, input_dim, hidden_dim, sigma_min, sigma_max, num_sigma):
3 super(ScoreNetwork, self).__init__()
4 [...]
5 self.sigmas = torch.exp(torch.linspace(
6 np.log(self.sigma_max), np.log(self.sigma_min), self.num_sigma))
7
8 # Neural network architecture
9 self.net = nn.Sequential(
10 nn.Linear(input_dim + 1, hidden_dim),
11 nn.ReLU(),
12 nn.Linear(hidden_dim, hidden_dim),
13 nn.ReLU(),
14 nn.Linear(hidden_dim, input_dim)
15 )
16
17 def forward(self, x, sigma):
18 # Concatenate sigma as an additional feature
19 sigma_feature = sigma.view(-1, 1)
20 x_sigma = torch.cat([x, sigma_feature], dim=1)
21 score = self.net(x_sigma)
22 return score
23
24 def loss_fn(self, x):
25 total_loss = 0.0
26 for sigma in self.sigmas:
27 sigma = sigma.to(device)
28 sigma = sigma.expand(x.size(0), 1)
29 noise = torch.randn_like(x) * sigma
30 x_noisy = x + noise
31 score = self.forward(x_noisy, sigma)
32 loss = ((score + noise / (sigma ** 2)) ** 2).mean() # Equation (4.6)
33 total_loss += loss
34 return total_loss
35
36 [...]
37 score_model = ScoreNetwork(input_dim, hidden_dim, sigma_min, sigma_max, num_sigma)
38
39 [...] # in training loop
40 loss = score_model.loss_fn(x) # Equation (4.6)
Input (x_noisy) Noise Level (sigma) Linear (4+1 $\to$ 512) ReLU Linear (512 $\to$ 512) ReLU Linear (512 $\to$ 4) Score (score)
Figure 13: Neural network structure of NCSN in HTS data
4.1.7   Results Comparison
Figure 14 presents the results of generating HTS data. In this section, we aim to share the empirical insights gained during the training of the models.
Similar to how VAEs tend to generate slightly blurry images in image generation, in this experiment with discrete values, the probability distribution also appeared relatively uniform rather than being overly skewed towards one side. However, it often struggled to accurately match the activity type and mode type, depending on the specified conditions. This may be attributed to the increased complexity resulting from the higher number of attributes.
In contrast, the GAN frequently encountered mode collapse, leading to cases where all probabilities were assigned to a particular activity or mode. However, this issue was less pronounced in cases involving fewer data attributes, such as problems related to origin location and destination location type.
The Normalizing Flow model demonstrated superior performance, with robustness across a wide range of hyperparameters. The loss reduction pattern followed the typical behavior expected in learning models, with an initial sharp decrease followed by a gradual decline in later stages. Throughout the training process and in the results, the model consistently performed well.
Figure 14: Results of each DGM in HTS data generation
The Diffusion model, which is commonly built using a U-Net architecture, was instead implemented with a simpler three-layer neural network in our approach. Despite initial concerns regarding its effectiveness, the model performed relatively well due to the simplicity of the dataset. However, occasional loss spikes occurred, which led the model to failure of training. Overall, the learning process remained stable, but future experiments could benefit from more sophisticated neural network architectures to further enhance performance.
The NCSN model exhibited high sensitivity to hyperparameters. For instance, when the max and min sigma values were low, the model displayed characteristics of high entropy or temperature, leading to unpredictable performance. Conversely, higher sigma values resulted in one-hot-like behavior. Additionally, the model’s performance varied across different categories: when it performed well on generating the location type, performance on generating types of action and mode became poor, and vice versa. Balancing hyperparameters effectively remains a key challenge for ensuring consistent performance across all attributes.
4.2 Generating Highway Traffic Speed Contour
The generation of highway traffic speed contours plays a crucial role in the analysis and visualization of traffic dynamics over both time and space. Traditionally, traffic flow has been represented using a time-space diagram of vehicles, which depicts the movement of vehicles along a road segment over time, providing insight into congestion patterns and traffic bottlenecks. However, converting this into a time-space speed contour offers several advantages. While traditional time-space diagrams focus on the position of vehicles at specific times, speed contours emphasize variations in traffic speed across both temporal and spatial dimensions, offering a clearer representation of traffic conditions such as slowdowns, jams, or free-flowing traffic across broader time periods. In this tutorial, we generated the traffic speed contour data based on the vehicle trajectory data.
Speed contours offer a continuous and intuitive visualization of traffic speeds across different locations and time intervals, making them particularly useful for traffic management and planning. They enable transportation professionals to quickly identify and quantify areas of concern, such as sections of highways prone to congestion or inconsistent traffic flow, thereby facilitating better decision-making. Moreover, these contours can serve as input for traffic forecasting models, aiding in the development of responsive traffic control systems, infrastructure improvements, and policy interventions aimed at reducing congestion and improving road safety.
4.2.1   Data and Preprocessing
Figure 15: Comparison of before and after the interpolation
In our tutorial section, we applied the DGMs to the Tennessee Department of Transportation’s I-24 Mobility Technology Interstate Observation Network (MOTION) data (Gloudemans et al., 2023) [^1] https://i24motion.org/data. I-24 MOTION data contains the vehicle trajectories in a four-mile section of I-24 in the Nashville-Davidson County Metropolitan area. It captures the trajectory in a high spatial and temporal resolution by using 294 ultra-high definition cameras and AI algorithms from Vanderbilt University. The AI algorithm transforms the raw videos into detailed 2-dimensional trajectories which is suitable to study vehicle behavior and traffic flow. The data collection occurred from Monday, November 21, 2022, to Friday, December 2, 2022, encompassing all eastbound and westbound lanes. The collection was done during peak morning hours, from 6 AM to 10 AM. This dataset is significant as it includes a variety of days as well as trajectories during an accident that occurred on November 21. However, it is important to note that when a single vehicle appears multiple times across different cameras, each instance is treated as a separate vehicle. Therefore, the whole trajectory will be sliced into multiple short sequences of trajectories. Consequently, this data needs a vehicle-matching algorithm to obtain the full trajectories of the single vehicle, but it can be effectively utilized for constructing time-space traffic contours without matching the vehicles.
The raw data from the MOTION project comprises segments of vehicle trajectory information. In this tutorial, our focus is on the reconstruction of traffic speed contours. To facilitate this, the trajectory data is transformed into velocity data corresponding to each specific time and position. Given that the calculation of velocity data based on trajectories inevitably results in the presence of zero or NaN values (except in the unlikely event of a vehicle being present at every possible position throughout the entire timespan), preprocessing was necessary. This preprocessing predominantly involved interpolation. The left and right of Figure 15 show the time-space traffic contours diagram before and after the interpolation.
The interpolation process is composed of two main steps. Firstly, we interpolate the zero and NaN values using linear and nearest-neighbor interpolation methods, respectively. Secondly, we apply a smoothing technique based on Edie’s definition of traffic flow dynamics (Treiber and Helbing, 2003). The initial interpolation step is crucial; without it, if we attempt to interpolate every zero or NaN value directly using Edie’s box method, the required size of Edie’s box would become excessively large, leading to significant distortion of the dataset. Before interpolation, most of the regions are colored in red because the 0 and NaN values are depicted as red. The actual values of these regions are much higher than what is shown in the figure. After interpolation, we can observe a much clearer propagation of congestion and its subsequent dissipation. Free flow is also clearly illustrated after interpolation. The blue region forming a straight line after 3,600 seconds indicates an increase in traffic speed following the morning rush hour as the time passes 9 AM.
Table 4: Pseudocode for data preprocessing
Algorithm 1: Smoothing the speed using the Edie’s definition
1 Given
2 $V(x,t)$: Velocity at $x,t$
3 $W,T$: Spatial and temporal width of the data
4 $\sigma$: Spatial range of smoothing
5 $\tau$: Temporal range of smoothing
6 $c_{free}$: Propagation velocity of perturbation in free flow
7 $c_{cong}$: Propagation velocity of perturbation in congested flow
8 $V_{c}$: Velocity that transition occurs from free flow to congested flow
9 $\Delta V$: Width of the transition region
10 for $x$00 to $W$, $t$00 to $T$ do
11 Initialize $\phi_{free}(x^{\prime},t^{\prime}),\phi_{cong}(x^{\prime},t^{\prime})$
12 For $x^{\prime}$$x-\nicefrac{{\sigma}}{{2}}$ to $x+\nicefrac{{\sigma}}{{2}}$, $t^{\prime}$$t-\nicefrac{{\tau}}{{2}}$ to $t+\nicefrac{{\tau}}{{2}}$ do
13 $\Delta x=x^{\prime}-x$
14 $\Delta t_{free}=(t^{\prime}-t)-\nicefrac{{(x^{\prime}-x)}}{{c_{free}}}$
15 $\Delta t_{cong}=(t^{\prime}-t)-\nicefrac{{(x^{\prime}-x)}}{{c_{cong}}}$
16 $\phi_{free}(x^{\prime},t^{\prime})=\exp(\nicefrac{{|\Delta x|}}{{\sigma}}-\nicefrac{{|\Delta t_{free}|}}{{\tau}})$
17 $\phi_{cong}(x^{\prime},t^{\prime})=\exp(\nicefrac{{|\Delta x|}}{{\sigma}}-\nicefrac{{|\Delta t_{cong}|}}{{\tau}})$
18 End for
19 Initialize $V_{free}(x,t),V_{free}(x,t)$
20 $V_{free}(x,t)=\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}(\phi_{free}(x^{\prime},t^{\prime})\times V(x^{\prime},t^{\prime})/\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\phi_{free}(x^{\prime},t^{\prime})$
21 $V_{cong}(x,t)=\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}(\phi_{cong}(x^{\prime},t^{\prime})\times V(x^{\prime},t^{\prime})/\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\phi_{cong}(x^{\prime},t^{\prime})$
22 $V_{min}(x,t)=min(V_{free}(x,t),V_{cong}(x,t))$
23 $tanh(x,t)=tanh(\nicefrac{{(V_{c}-V_{min}(x,t))}}{{\Delta V}})$
24 $w=0.5\times(1+tahnh(x,t)$
25 $V(x,t)=(1-w)\times V_{free}(x,t)+w\times V_{cong}(x,t)$
26 End for
The entire process of interpolation, which is based on Edie’s definition of traffic flow dynamics, is outlined in Table 4. The smoothing process, known as the “adaptive smoothing method,” is designed to filter out noisy fluctuations while considering the primary direction of flow propagation. This method employs spatiotemporal correlation analysis to identify the dominant characteristic line. The adaptive smoothing method uses a spatio-temporal low-pass filter that allows only low-frequency Fourier components to pass through, smoothing out high-frequency components. The filter eliminates high-frequency noises over a timescale shorter than $\tau$ and spatial noise over a length scale shorter than $\sigma$. The values for $\tau$ and $\sigma$ were determined through empirical trials to ensure the data is not overly blurred while effectively passing the main propagation. The filter captures two main propagation types: free flow and congested traffic flow. These two propagations have different coefficients of velocity and direction, reflecting the distinct properties of each traffic flow type. The property of each traffic flow is reflected based on the different values of each coefficient.
4.2.2   Variational Autoencoder
We utilize the result of Equation (4.5) to compute the loss of the VAE. The code structure and implementation of the loss function are consistent with the approach presented in Section 4.1.2. Figure 16 illustrates the structure of the neural network for traffic speed contour generation. Figure 16 (a) illustrates the structure of the encoder. The encoder encodes the data to Mean and Variance with a $D$-dimensional vector. In this section, $D=64$ is applied. Figure 16 (a) is the structure of the decoder that generates the image from the latent vector.
Figure 16: Neural network structure of VAE in highway traffic speed contour
Figure 17: Neural network structure of GAN in highway traffic speed contour
4.2.3   Generative Adversarial Networks
The loss function and model architecture of the GAN for traffic speed contour estimation follows the implementation outlined in Section 4.1.3. We employ binary cross-entropy loss for both the discriminator and the generator.
Figure 17 presents the detailed neural network architecture of the GAN, which closely resembles the VAE structure depicted in Figure 16. The discriminator shown in Figure 17 (a) outputs the decision for the input if it is real or fake. The generator, shown in Figure 17 (b) outputs the image from the $D$-dimensional noise vector. Here, $D=64$ is applied. This similarity of the neural network structure between GAN and VAE enables us to qualitatively compare the performance of these two models. The models discussed in subsequent sections, such as Normalizing Flows, Score-based Generative Models, and Diffusion Models, possess more intricate and distinctive structures. Consequently, evaluating these models within the same neural network framework is impractical. However, the architectures of the VAE and GAN can be designed to be similar, which fulfills our purpose of comparison.
4.2.4   Normalizing Flows (Flow-based Generative Models)
As we discussed at Section 2.5, among various Flow-based Generative Models such as Dinh et al. (2014); Rezende and Mohamed (2015); Kingma and Dhariwal (2018); Durkan et al. (2019), we implemented RealNVP (Dinh et al., 2016) because of its ability to generate high-dimensional distribution due to its flexibility. As we can see in Equation (2.21), we need to calculate the $\log p_{0}(z_{0})$ in Equation (2.20) to get the likelihood of the data. In the implementation of RealNVP, we can simplify the log-likelihood $\log\left(p(\mathbf{z})\right)$ by assuming $z$ is sampled from the standard normal distribution. The other term in Equation (2.20), the summation of the determinant of the Jacobian, is simplified by applying a coupling layer as described through Equation (2.22) to Equation (2.26). Thus, the summation of the determinant of the Jacobian can be simplified as $\sum_{j}s(x_{1:d})_{j}$. In our code application, the regularization loss for batch normalization and scale parameters are applied following the contents in RealNVP (Dinh et al., 2016).
Dinh et al. (2016) multiplied $5\times 10^{-5}$ for the regularization term of scale network. However, since the coefficients must be re-calibrated based on the data, we fine-tuned each coefficient for the loss term in our implementation. A general guideline is to prioritize log-likelihood over the sum of the determinant of the Jacobian. Additionally, the sum of the batch normalization also plays a critical role in the qualitative performance, which has been incorporated into the provided code.
Figure 18: Neural network structure of RealNVP in highway traffic speed contour
The neural network structure of RealNVP is shown in Figure 18. The neural network is composed of multiple layers of its Backbone Block, which is shown in Figure 18 (a). Each Backbone Block consists of convolution, Instance 2D Normalization, and ReLU layer. Each Backbone Block is the smallest unit for the scale and translation network. In the implementation in the current paper, we stacked 22 layers of scale and translation network. The dimensions and number of channels of the features change at each layer, as detailed in Figure 18 (b). In the generation stage, the computation proceeds in an inverse way.
Figure 19: Neural network structure of DDPM in highway traffic speed contour
4.2.5   Diffusion Models
The loss function of DDPM is simplified as the difference between the predicted noise and the real noise as stated in Equation (2.39). The criteria of the difference, i.e., the norm of difference, can be varied by the problem. Our code is also designed to select one of three types of loss, which is $L_{1}$, Mean Square Error (MSE) loss, and Huber loss. For simplicity, we inserted the code block only including the MSE loss.
The neural network that is used for the DDPM is illustrated in Figure 19. Figure 19 (a), Figure 19 (b), and Figure 19 (c) shows the structure of the Residual Block, Backbone Block, and U-Net of DDPM respectively. The U-Net structure that is applied in the neural network in DDPM enables one to learn the information from high-level features or features from previous layers. The residual block contains the time embedding as an input, which identifies the stage of the current generation process. A similar mechanism can also be found in the Score-based model, as the score-based model creates an image using the scheduled $\sigma$. The neural network in DDPM also includes the attention mechanism that captures the relationship between features. The attention mechanism was applied inside the Backbone block, between the Residual block and the group normalization.
4.2.6   Score-based Generative Models
In the code implementation of NCSN in traffic speed contour, we had minor adjustments from the code from Section 4.1.6. By definition, $\tilde{\mathbf{x}}$ is defined as $\tilde{\mathbf{x}}=\mathbf{x}+\sigma\cdot\epsilon$. Then $\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)$ in Equation (4.7) is transformed as $\nicefrac{{\epsilon}}{{\sigma}}$. Practically, $\sigma^{2}$ is multiplied by the objective function as the weight of each stage. The detailed implementation can be found in the code.
The neural network of NCSN is formed with a block-wise structure with a Residual Block. Figure 20 (a) and Figure 20 (b) illustrate the structure of the Residual Block and whole neural network for score function, respectively. Skip-connection is applied in each block to maintain the low-level feature information.
Figure 20: Neural network structure of NCSN in highway traffic speed contour
Figure 21: Results of each DGM in highway traffic speed contour generation
4.2.7   Results Comparison
Figure 21 presents the outputs generated by various models, including ground truth data, VAE, GAN, RealNVP, DDPM, and NCSN. It is evident that the VAE outputs are noticeably more blurred compared to those produced by the other models. Throughout the training process, we experimented with a wide range of neural network architectures for the VAE. Despite attempting to train the VAE with neural networks that had four to eight times more parameters than the current configuration, we were unable to eliminate the inherent blurriness of the VAE outputs. Unlike other algorithms, which exhibited some variability in performance across training sessions, the VAE consistently produced stable results both numerically and visually. However, the VAE outputs remained uniformly blurred, with some generated images appearing severely noisy and closely resembling the mean of the dataset.
In contrast, GANs demonstrated superior performance, especially given the simplicity of their neural network architecture. Even when using a neural network architecture almost identical to that of the VAE, the GAN produced clearer and more diverse images. Although the training process occasionally showed abrupt fluctuations in the loss function, as long as the generator’s training did not lag significantly behind that of the discriminator, the model eventually achieved a balance. Considering the simplicity of the code, training time, and generation speed, the strong performance of such a straightforward algorithm is very encouraging.
RealNVP displayed the most variability among all the algorithms. Occasionally, it produced excellent results, but this occurred in only about 1 in 20 training runs. Since the loss function consists of four terms, the model’s training trajectory varied depending on the coefficients applied to each term. The coefficients we proposed in the code were the result of extensive trial and error. It was observed that both log-likelihood and regularization were crucial, and the performance was highly sensitive to the numerical value of the first term in each component. Additionally, once RealNVP fell into a local minimum during training, it was unable to escape, resulting in poor performance across all samples and mode collapse, where the images appeared more like homogeneous noise. The generated images resembled ground truth data with slightly mixed boundaries between data values.
Compared to RealNVP, DDPM demonstrated relatively stable performance. It consistently produced uniform models with good image quality across different training sessions. The resulting images resembled ground truth data with the addition of white noise, similar to photographs taken at a very high ISO setting. NCSN, which follows a mechanism fundamentally similar to that of DDPM, exhibited comparable performance. Despite the uniform resolution of 64x64, the images conveyed a sense of being highly fine-grained.
However, it is crucial to approach model performance evaluations with caution. Qualitative comparisons do not provide a comprehensive assessment of model efficacy. The performance of these models can vary significantly based on hyperparameter tuning and architectural choices, making it challenging to definitively claim the superiority of any single neural network configuration or hyperparameter set. Interestingly, our empirical observations during implementation revealed instances where simpler network structures yielded superior results, underscoring the complexity of model optimization. This phenomenon highlights the importance of thorough experimentation and the potential pitfalls of overgeneralization in neural network performance assessment. Future research should focus on developing more robust quantitative metrics for model comparison, considering the multifaceted nature of generative model performance across various tasks and datasets.
Challenges and Opportunities
5.1 Evaluation and Validation of Generative Models
In transportation research, evaluating the performance of DGMs poses several challenges, particularly in unsupervised learning applications such as data imputation, anomaly detection, and synthetic data generation, due to the high-dimensional nature of the underlying probability distribution. In traffic data imputation, DGMs are used to fill in missing values in sensor data, such as loop detector data or GPS vehicle trajectories. Given that the ground truth values for missing data are often unknown, assessing how well the imputed data captures the true underlying traffic patterns is challenging, especially considering the complex spatiotemporal dependencies in traffic data. This often requires a carefully designed cross-validation scheme to ensure the credibility of DGMs. In anomaly detection, DGMs are trained to learn the underlying distribution of normal data in an unsupervised manner, allowing them to identify deviations or anomalies that are outside the learned normal data distribution. However, the causes of anomalies are often not unknown a priori, and defining what constitutes normal and anomalous behaviors in traffic data can be subjective and context-dependent. As such, preparing training data that contain only normal instances for DGMs is a challenge. In data generation, DGMs are employed to generate synthetic data for tasks such as data augmentation, traffic simulation, and privacy-preserving data sharing. However, assessing the realism and utility of the generated synthetic traffic data is challenging due to the absence of ground truth and the subjective nature of realism. For example, in the generation of synthetic vehicle trajectory, there is a lack of standardized evaluation metrics that comprehensively measure the similarity between generated and true population trajectory datasets, which encompass critical aspects such as route distribution, origin-destination distribution, travel time distributions, and individual driving patterns.
Evaluating DGMs involves assessing how well the models is able to generate data that match the underlying distribution of the observed data. A critical limitation in transportation data analysis is that the amount of available data is small and deep learning models could overfit to the training data without proper evaluation and validation. In this context, scoring rules provide a principled way to evaluate these models by quantifying how well the predicted probabilities align with the actual outcomes (Gneiting and Raftery, 2007). For instance, the Log Score measures the negative log-likelihood of the generative data under the probabilistic distribution derived by DGMs. The Continuously Ranked Probabilistic Score (CRPS) and Energy Score (ES) measure how well the cumulative distribution obtained produces continuous univariate and multivariate data, respectively. A second challenge is that many types of traffic data are governed by “physics”. For example, the probability distribution of the trajectory of a car needs to be governed by the continuous variation of acceleration and turning angle. If one only considers location information in building the DGMs, the trained model will hardly reproduce realistic trajectories, which poses further problems in applying DGMs as data generators for downstream tasks. Another example is the traffic speed data that we introduced in Sec 4.2. The speed in this case is produced by microscopic traffic flow. Treating these data as an image will lose the physical nature behind it, and unrealistic data that violate traffic flow patterns could be generated. A third challenge is how to define a proper probability distribution for complex structured data. Taking the travel survey data presented in Sec 4.1 as an example, multiple constraints exist at different hierarchical levels (see e.g., Sun et al. (2018)): the composition of a household will be governed by certain structures; individuals within the same household exhibit strong interactions/correlation in term of social demographic features; trips of an individual have to satisfy certain sequence orders and in the meanwhile trips chains of different individuals will also be coupled (e.g., a trip of an adult to drop off a child at school). To model such data, one often has to introduce strong but unrealistic assumptions to simplify the probabilistic model, such as assuming household members are independent of each other. These assumptions make the computational part easier, but in the meanwhile the quality of data will be undermined.
5.2 Dynamic Characteristics of Transportation data
Mobility patterns and traffic data are continually evolving due to factors such as seasonal variations, changes in travel behavior, infrastructure updates, and technological advancements. These evolving conditions can lead to shifts in the underlying data distribution, posing challenges for DGMs trained on historical data. Such shifts can decrease the performance and reliability of the trained DGMs, especially in long-term deployments or applications where transportation patterns change over time. For instance, in tasks like driving scenario generation or population synthesis, the introduction of new vehicle types (e.g., electric or autonomous vehicles) or emerging transportation modes (e.g., shared micro-mobility) can alter driving behaviors and mode choice patterns, making the synthetic datasets generated by DGMs outdated. In anomaly data generation tasks, where anomalies are outliers in the data distribution (e.g., incidents, road blockages, or sudden traffic surges), dynamic changes in transportation can cause shifts in both normal and anomaly data distributions. This makes it harder to detect shifts in anomaly patterns and further complicates the distinction between normal and anomalous data. In estimation and prediction tasks, distribution shifts at different levels pose their own challenges. At the agent level, shifts in pedestrian trajectory data may occur due to behavioral changes driven by public health events, such as the COVID-19 pandemic. During such events, pedestrians tend to maintain larger physical distances, altering typical movement patterns and interactions in public spaces. Similarly, vehicle trajectory distributions can shift due to the introduction of autonomous vehicles or vehicle-to-vehicle communications that affect the way vehicles interact with traffic and can alter driving patterns. At the link level, new road construction or traffic control strategies can change the flow, speed, or density of links, making models trained on previous data less accurate. At the regional level, gradual shifts in population demographics, land use, or public transit services can alter travel demand, leading to outdated models for Origin-Destination (OD) estimation or demand forecasting.
To enable DGMs to effectively manage and adapt to data distribution shifts in evolving transportation environments, strategies can be approached in two key steps: detecting changes in data distribution and updating the model to incorporate new datasets. The first step, detecting changes in data distribution, involves identifying when and how the characteristics of the data have evolved from what was seen during the model’s training phase. This can be achieved using statistical tests, such as the Kolmogorov-Smirnov test for numerical features or the Chi-square test for categorical features, which assess whether significant differences exist between the distribution of new data and the original training dataset. However, these tests are typically used for single or pairs of features, and applying these tests to high-dimensional datasets can be computationally expensive and may not effectively capture interactions between features. In high-dimensional datasets, distance metrics like Wasserstein Distance, Jensen-Shannon Divergence, Kullback-Leibler Divergence, or Cosine Similarity can be used to quantitatively measure how far apart the distributions are. Additionally, visualization techniques such as histogram analysis and dimensionality reduction methods like t-distributed Stochastic Neighbor Embedding (t-SNE) or Principal Component Analysis (PCA) can provide visual insights into changes in feature distributions. DGMs themselves can also be leveraged to detect changes in data distributions by comparing the distributions of new data to those of the training data using the DGM’s data generation and density estimation capabilities. Once a shift in data distribution has been detected, the next step is to update the model to incorporate new trends. The model may be retrained at regular intervals (periodic retraining) or in response to specific triggers or thresholds (trigger-based retraining). However, retraining the model from scratch with the new data can be computationally intensive and time-consuming, particularly for large datasets and complex models. An alternative approach is to employ incremental learning or continual learning methods, which update the model progressively as new data arrives, avoiding the need to retrain the entire dataset. Fine-tuning through transfer learning is another strategy, where a pre-trained model is adjusted to better align with the new distribution while retaining previously learned features. Fine-tuning involves updating a subset of the model parameters using a small amount of labeled data from the new distribution, providing a computationally efficient means to adapt to evolving data.
5.3 Deep Generative Models for Enhancing Transportation Data Privacy
Synthetic data generation is one of the core applications of DGMs, as discussed in Section 3.1.1. The use of synthetic data generated by DGMs offers several significant advantages for privacy protection in transportation, highlighting the potential of synthetic data to balance the need for data utility with stringent privacy requirements. One of the primary benefits of synthetic data generated by DGMs is its ability to preserve the statistical or distributional properties and patterns of real-world transportation and mobility data while excluding personally identifiable information. This means that the generated synthetic data can reflect the trends, behaviors, and characteristics found in actual data without risking individual privacy. For instance, in trajectory data containing personal origin-destination pairs and commuting patterns (Choi et al., 2018), synthetic data can accurately represent peak travel times, popular routes, and average journey durations without including any specific details about individual travelers. This capability ensures that the generated synthetic data is useful for analysis, model training, and validation. Researchers and practitioners can derive meaningful insights for transportation policy and develop effective models for traffic management without compromising the quality and accuracy of the data. For example, urban planners can use synthetic data to simulate the impact of new infrastructure projects on traffic flow, while data scientists can train machine learning models to predict congestion or optimize public transport schedules. Furthermore, synthetic data enables safe data sharing and collaboration between different organizations, stakeholders, and researchers. For instance, transportation agencies who collected privacy-sensitive trajectory data, activity sequence, and time use data, can share synthetic datasets generated by DGMs with third-party developers, policymakers, and academic institutions without exposing sensitive information. This can foster innovation, accelerate research, and support the development of new solutions and services, all while maintaining privacy.
5.4 Addressing Data Bias for Equitable Data Generation
The performance of DGMs heavily depends on the quality and representativeness of the training data. Ideally, Generative AIs can generate realistic and unbiased data if the training data itself is unbiased. However, in practice, training data often contains inherent biases, whether related to socioeconomic factors, geographic disparities, or historical inequalities. For example, if a generative model for urban planning is trained on data predominantly from affluent neighborhoods, the synthetic data it produces might overemphasize characteristics specific to those areas, neglecting the needs and patterns of lower-income regions. This could result in urban planning decisions that disproportionately benefit wealthier communities, thereby perpetuating or even amplifying existing disparities. As a result, ensuring that generative models do not amplify these biases is crucial for producing fair and representative synthetic data. This requires a multifaceted approach, beginning with the careful selection and preprocessing of training data. Data preprocessing steps might include normalizing data distributions, balancing underrepresented groups, and removing explicit bias indicators. For instance, transportation data, it might involve ensuring that data from various demographics and geographic regions are equally represented.
5.5 Towards Trustworthy AI
Beyond implementing technical measures, securing public and stakeholder acceptance and building trust in synthetic data presents significant challenges. As the use of synthetic data becomes more prevalent in AI systems, ensuring that these systems embody trustworthy AI principles is essential. Concerns may arise about the reliability and credibility of policies and decisions based on synthetic data. To promote trustworthy AI, it is crucial to maintain transparency in the documentation of data generation processes, open-sourcing code for reproducibility, and explaining the methodologies and safeguards used making them accessible to non-technical stakeholders. Additionally, demonstrating the validity of synthetic data through rigorous testing and comparison with real-world outcomes reinforces trustworthiness. For example, when using synthetic traffic data for new policy design, models should be able to replicate real traffic patterns under various conditions to demonstrate their adaptability and reliability. Fairness in synthetic data generation is equally important, as DGMs must be designed to avoid perpetuating biases inherent in training data. Employing fairness-aware algorithms and conducting regular audits of the synthetic data can help identify and mitigate biases, ensuring ethical and equitable outcomes. Moreover, involving stakeholders—such as community representatives and subject matter experts—in the development and implementation process further enhances credibility and acceptance. Their insights can help identify potential biases in data and modeling, as well as practical implications to ensure that synthetic data applications align with societal needs and ethical standards, thereby fostering trustworthy AI.
5.6 Towards Causal Generative Models
Causality refers to the relationship between cause and effect, where one event (the cause) directly influences another event (the effect). Assuming that event $\mathbf{x}$ is the direct cause of event $\mathbf{y}$, the causal relationship $\mathbf{x}\rightarrow\mathbf{y}$ signifies that altering $X$ can lead to a change in $\mathbf{y}$, whereas the reverse cannot hold true (Pearl, 1995, 2009). The concept is fundamental in understanding how and why certain phenomena occur, especially in complex systems such as transportation networks. Different from traditional methods that rely on statistical inference, which usually calculates the conditional probability $p(\mathbf{y}|\mathbf{x})$ and cannot consider confounding bias, causal models focus on uncovering the true cause-and-effect relationships. In other words, these models can isolate the impact of specific variables and provide a deeper understanding of how changes in $\mathbf{x}$ directly lead to changes in $\mathbf{y}$ (Wagner, 1999; Pearl and Bareinboim, 2014; Bagi et al., 2023). In transportation research, causality has been actively considered in many areas such as traffic flow (Queen and Albers, 2009; Li et al., 2015; Du et al., 2023), urban planning (Huang et al., 2022a; Akbari et al., 2023), and traffic safety (Davis, 2000; Elvik, 2011; Mannering et al., 2020) to understand the impact of various interventions on outcomes. Studies in these areas evaluate the effects of variables that contribute to issues like traffic bottlenecks, vulnerabilities, and severe accidents. These applications demonstrate the value of causal inference in uncovering the true triggers of transportation phenomena, leading to more effective interventions and policy decisions.
In transportation applications, integrating causality into DGMs can significantly enhance their effectiveness. Conventional generative models primarily rely on training data, which limits their effectiveness in new or changing environments. In contrast, causal generative models leverage underlying cause-effect relationships, enhancing their ability to generalize across different settings and maintain robustness in out-of-distribution (OOD) scenarios. By learning causality, these models can adapt to covariate shifts—where input data distributions vary between training and deployment—with fewer samples, reusing many of their components without the need for retraining. For example, a causal generative model trained to understand the relationship between traffic density and travel time in one city can adapt to another city with different traffic patterns. Additionally, causal generative models maintain reliable predictions under OOD conditions such as major social events, natural disasters, or unexpected road closures by understanding and leveraging causal structures. This dual capability of improved generalization and OOD robustness makes causal generative models particularly valuable for developing scalable and effective transportation solutions in diverse and dynamically changing environments, as discussed in Section 5.2. Moreover, considering causality can improve the interpretability of DGMs. This capability not only helps stakeholders understand and trust the model’s outputs but also allows them to control the generative process to produce targeted data. Traditional deep learning models are often perceived as black boxes, providing limited insights into the reasons behind certain outputs. Conversely, causal generative models focus on causal mechanisms, which offer clearer explanations for transportation phenomena. For example, in the context of traffic safety, a causal generative model can not only predict the likelihood of accidents at a particular intersection but also explain contributing factors to output such as poor road design, inadequate road signs, or high pedestrian activity. This interpretability is essential for developing trustworthy AI, as discussed in Section 5.5 Since the model is interpretable based on these causal factors, it also allows for the control of the generative process by directly manipulating these variables. This kind of controllable generative model, grounded in causality, can be used to generate data representing specific scenarios, enabling transportation planners to simulate and evaluate the effects of interventions. This combination of interpretability and controllability makes causal generative models highly valuable in practical transportation applications.
Conclusion
This review paper offers a comprehensive review of state-of-the-art DGMs, which are transforming how to understand the underlying patterns of increasingly large amount of data. This paper also offers a review of state-of-art research in various transportation-related topics using DGMs, examining the current landscape and practices in transportation research with DGMs. Furthermore, this paper includes an open-sourced tutorial, offering practical guidance and resources for those looking to explore and implement DGMs in transportation research and beyond. Finally, this paper offers discussions on potential challenges and opportunities related to utilizing, implementing, and developing DGMs in the transportation domain. The authors hope that this paper contributes to the advancement of the transportation academic field and promotes further research in DGMs.
[…]
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT and Grammarly for language assistance. After utilizing these tools, the author(s) thoroughly reviewed and revised the content as necessary and take full responsibility for the final version of the published article.
Data and Code Availability Statement
All data and code associated with this work are available at https://github.com/benchoi93/DGMinTransportation.
Appendix A Derivation for Diffusion Model
Similar to VAE and Normalizing Flows (models using explicit densities for training), the training objective is to maximize the likelihood, or minimize the negative log-likelihood as follows:
$\theta^{*}=\arg\min_{\theta}\mathbbm{E}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right]$ (A.1)
Similar to VAE, instead of directly training the model with negative log-likelihood, we can consider $q$ as the approximate posterior and use the variational bound on negative log-likelihood to train the model as follows:
$\begin{split}&\mathbbm{E}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right],\\ &=\mathbbm{E}\left[-\log\frac{p_{\theta}\left(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T}\right)}{p_{\theta}\left(\mathbf{x}_{1},\cdots,\mathbf{x}_{T}|\mathbf{x}_{0}\right)}\right],&&\Rightarrow\text{Baye's Rule}\\ &=\mathbbm{E}\left[-\log\frac{p_{\theta}\left(\mathbf{x}_{0:T}\right)}{p_{\theta}\left(\mathbf{x}_{1:T}|\mathbf{x}_{0}\right)}\cdot\frac{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],\\ &=\mathbbm{E}\left[-\log\frac{p_{\theta}\left(\mathbf{x}_{0:T}\right)}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\cdot\frac{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}{p_{\theta}\left(\mathbf{x}_{1:T}|\mathbf{x}_{0}\right)}\right]=\mathbbm{E}\left[-\log\frac{p_{\theta}\left(\mathbf{x}_{0:T}\right)}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right]-\mathbbm{E}\left[\log\frac{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}{p_{\theta}\left(\mathbf{x}_{1:T}|\mathbf{x}_{0}\right)}\right],\\ &\leq\mathbbm{E}\left[-\log\frac{p_{\theta}\left(\mathbf{x}_{0:T}\right)}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],&&\Rightarrow\text{KL divergence }\geq 0\\ \end{split}$ (A.2)
Using Equation (2.34) and Equation (2.35), we can further derive:
$\begin{split}&\mathbbm{E}\left[-\log\frac{p_{\theta}\left(\mathbf{x}_{0:T}\right)}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],\\ &=\mathbbm{E}\left[-\log\frac{p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{\prod_{t=1}^{T}q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)}\right],\\ &=\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\log\frac{\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{\prod_{t=1}^{T}q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)}\right],\\ &=\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=1}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)}\right],\\ \end{split}$ (A.3)
Appendix A in Ho et al. (2020) further provides a detailed derivation of the reduced variance variational bound for diffusion models as follows:
$\begin{split}&\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=1}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)}\right],\\ &=\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)}-\log\frac{p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)}\right],\\ &=\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)}\cdot\frac{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)}-\log\frac{p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)}\right],\\ \end{split}$ (A.4)
since
$\begin{split}q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1}\right)&=q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{x}_{0}\right),&&\Rightarrow\text{Markov chain property}\\ &=\frac{q\left(\mathbf{x}_{t},\mathbf{x}_{t-1},\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t-1},\mathbf{x}_{0}\right)}=\frac{q\left(\mathbf{x}_{t},\mathbf{x}_{t-1},\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t-1},\mathbf{x}_{0}\right)}\cdot\frac{q\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)},\\ &=\frac{q\left(\mathbf{x}_{t},\mathbf{x}_{t-1},\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)}\cdot\frac{q\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t-1},\mathbf{x}_{0}\right)},\\ &=q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)\cdot\frac{q\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t-1},\mathbf{x}_{0}\right)}.\\ \end{split}$ (A.5)
As a result,
$\begin{split}&\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)}\cdot\frac{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)}-\log\frac{p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)}\right],\\ &=\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)}-\sum_{t=2}^{T}\log\frac{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)}-\log\frac{p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)}\right],\\ &=\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)}+\log q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)-\log p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)\right],\\ \end{split}$ (A.6)
since
$\begin{split}&-\sum_{t=2}^{T}\log\frac{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)}-\log\frac{p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)},\\ &=-\log\frac{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{2}|\mathbf{x}_{0}\right)}\cdot\frac{q\left(\mathbf{x}_{2}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{3}|\mathbf{x}_{0}\right)}\cdot\frac{q\left(\mathbf{x}_{3}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{4}|\mathbf{x}_{0}\right)}\cdots\frac{q\left(\mathbf{x}_{T-1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)}\cdot\frac{p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}{q\left(\mathbf{x}_{1}|\mathbf{x}_{0}\right)},\\ &=\log q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)-\log p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right).\\ \end{split}$ (A.7)
Therefore,
$\begin{split}&\mathbbm{E}\left[-\log p\left(\mathbf{x}_{T}\right)-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)}+\log q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)-\log p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)\right],\\ &=\mathbbm{E}\left[\log\frac{q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)}{p\left(\mathbf{x}_{T}\right)}-\sum_{t=2}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)}-\log p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)\right],\\ &=\mathbbm{E}\left[\underbrace{D_{KL}\left(q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)||p\left(\mathbf{x}_{T}\right)\right)}_{L_{T}}+\sum_{t=2}^{T}\underbrace{D_{KL}\left(q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)||p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)\right)}_{L_{t-1}}\underbrace{-\log p_{\theta}\left(\mathbf{x}_{0}|\mathbf{x}_{1}\right)}_{L_{0}}\right].\\ \end{split}$ (A.8)
Therefore, the overall loss function of minimizing the negative log-likelihood in Equation (A.1) is decomposed into several losses, $L_{T}$, $L_{t-1}$, and $L_{0}$. Here, $L_{T}$ is constant since both $q\left(\mathbf{x}_{T}|\mathbf{x}_{0}\right)$ and $p\left(\mathbf{x}_{T}\right)$ are fixed, and therefore, we can ignore this term. Also, in Ho and Ermon (2016), $L_{0}$ is explicitly defined by using the characteristics of the image generation problem, and as a result, $L_{0}$ can be interpreted as a reconstruction loss of a problem-specific decoder. As a result, the actual learning process of the diffusion model is related to $L_{t-1}$.
$L_{t-1}$ measures the KL-divergence of $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$ from $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$. The diffusion process, $q$, represents the process of adding small noise to the data; i.e., given a less noisy data $\mathbf{x}_{t-1}$, the distribution of a more noisy data $\mathbf{x}_{t}$. The first term, $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$, represents the true denoising process which is derived from the definition of $q$ given the true data without noise, $\mathbf{x}_{0}$. What the diffusion models try to learn is the denoising process $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$; i.e., given a more noisy data $\mathbf{x}_{t}$, the distribution of a less noisy data $\mathbf{x}_{t-1}$. As a result, $L_{t-1}$ captures the distributional difference between the true denoising process (given the true data) and the approximated denoising process (without the true data).
Since the diffusion process follows Gaussian distribution, the true reverse process, $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$, can be assumed to follow a Gaussian distribution if $T$ is sufficiently large, or $T\rightarrow\infty$. Let $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)=\mathcal{N}(\mathbf{x}_{t-1};\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0}),\tilde{\beta}_{t}\mathbbm{I})$. To derive explicit form of $\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0})$ and $\tilde{\beta}_{t}$, first, we should derive a closed-form equation for sampling $\mathbf{x}_{t}$ at an arbitrary timestep $t$ from Equation (2.32) as follows:
$\begin{split}\mathbf{x}_{t}&=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\sqrt{\beta_{t}}\boldsymbol{\epsilon}_{t},&&\\ &=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t},&&\Rightarrow\text{Let }\alpha_{t}=1-\beta_{t}\\ &=\sqrt{\alpha_{t}}\left(\sqrt{\alpha_{t-1}}\mathbf{x}_{t-2}+\sqrt{1-\alpha_{t-1}}\boldsymbol{\epsilon}_{t-2}\right)+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t},&&\\ &=\sqrt{\alpha_{t}\alpha_{t-1}}\mathbf{x}_{t-2}+\sqrt{1-\alpha_{t}\alpha_{t-1}}\bar{\boldsymbol{\epsilon}}_{t},&&\Rightarrow\bar{\boldsymbol{\epsilon}}_{t}\text{ merges two Gaussians }(\boldsymbol{\epsilon}_{t}\text{ and }\boldsymbol{\epsilon}_{t-1})\\ &=\cdots\\ &=\sqrt{\alpha_{t}\alpha_{t-1}\cdots\alpha_{1}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}\alpha_{t-1}\cdots\alpha_{1}}\boldsymbol{\epsilon},&&\\ &=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},&&\Rightarrow\text{reparameterize }\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}\\ \end{split}$ (A.9)
therefore,
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\tilde{\alpha}}_{t}\mathbf{x}_{0},(1-\tilde{\alpha_{t}})\mathbbm{I}).$ (A.10)
Then,
$\begin{split}q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)&=q\left(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{x}_{0}\right)\frac{q\left(\mathbf{x}_{t-1}|\mathbf{x}_{0}\right)}{q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)},\\ &\propto\exp\left(-\frac{1}{2}\left(\frac{(\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{t-1})^{2}}{\beta_{t}}+\frac{(\mathbf{x}_{t-1}-\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0})^{2}}{1-\bar{\alpha}_{t-1}}-\frac{(\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0})^{2}}{1-\bar{\alpha}_{t}}\right)\right),\\ &=\exp\left(-\frac{1}{2}\left(\left(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}\right)\mathbf{x}_{t-1}^{2}-\left(\frac{2\sqrt{\alpha_{t}}\mathbf{x}_{t}}{\beta_{t}}+\frac{2\sqrt{\bar{\alpha}_{t-1}\mathbf{x}_{0}}}{1-\bar{\alpha}_{t-1}}\right)\mathbf{x}_{t-1}+\cdots\right)\right),\\ \end{split}$ (A.11)
where $\cdots$ include the terms irrelevant to $\mathbf{x}_{t-1}$. Since $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$ is a Gaussian distribution, we can find the mean and the variance from Equation (A.11):
$\begin{split}&\tilde{\beta}_{t}=\frac{1}{\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t},\\ \end{split}$ (A.12)
and
$\begin{split}\tilde{\boldsymbol{\mu}}_{t}&=\left(\frac{\sqrt{\alpha_{t}}\mathbf{x}_{t}}{\beta_{t}}+\frac{\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}}{1-\bar{\alpha}_{t-1}}\right)/\left(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}\right),\\ &=\left(\frac{\sqrt{\alpha_{t}}\mathbf{x}_{t}}{\beta_{t}}+\frac{\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}}{1-\bar{\alpha}_{t-1}}\right)\cdot\left(\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}\right),\\ &=\frac{(1-\bar{\alpha}_{t-1})\sqrt{\alpha_{t}}}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\mathbf{x}_{0},\\ &=\frac{(1-\bar{\alpha}_{t-1})\sqrt{\alpha_{t}}}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\left(\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t})\right),\\ &=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{t}\right).\\ \end{split}$ (A.13)
Therefore, we can rewrite the $L_{t-1}$ term in Equation (A.8) using the KL-divergence between two Gaussian distributions as:
$L_{t-1}=D_{KL}\left(q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)||p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)\right)=\mathbbm{E}_{q}\left[\frac{1}{2\sigma_{t}^{2}}\|\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0})-\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right]+C.$ (A.14)
Essentially, $\mu_{\theta}(\mathbf{x}_{t},t)$ is a $\theta$-parameterized function that predicts the distribution mean given $\mathbf{x}_{t}$ and $t$. We can reparameterize it as:
$\begin{split}\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right),\end{split}$ (A.15)
which then
$\begin{split}L_{t-1}&=\mathbbm{E}_{q}\left[\frac{1}{2\sigma_{t}^{2}}\left\lVert\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0})-\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\right\rVert^{2}\right]+C,\\ &=\mathbbm{E}_{q}\left[\frac{(1-\alpha_{t})^{2}}{2\alpha_{t}(1-\alpha_{t})\sigma_{t}^{2}}\left\lVert\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right\rVert^{2}\right]+C,\\ &=\mathbbm{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\frac{(1-\alpha_{t})^{2}}{2\alpha_{t}(1-\alpha_{t})\sigma_{t}^{2}}\left\lVert\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}}_{t}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},t\right)\right\rVert^{2}\right]+C.\\ \end{split}$ (A.16)
Appendix B Derivation for Score-based Generative Model
Once the score network is trained, data generation can be accomplished through an iterative method known as Langevin dynamics, or Langevin Monte Carlo (Parisi, 1981; Grenander and Miller, 1994). Langevin dynamics is originally formulated to describe the behavior of molecular systems, such as Brownian motion. The fundamental principle of Langevin dynamics in the context of Brownian motion is captured by the Langevin Equation, which is expressed as follows:
$\begin{split}m\ddot{x}=-\lambda\dot{x}+\eta\end{split},$ (B.1)
where $m$ is the mass of the particle, $\lambda$ is the damping coefficient, and $\eta\sim\mathcal{N}(0,2\sigma^{2})$ is a Gaussian noise.
The dynamics of Brownian motion as described in Equation (B.1) can be further derived by considering the potential energy in the system. Specifically, the force acting on a particle can be expressed in terms of the gradient of potential energy; i.e., $\frac{\partial V(x)}{\partial x}=\nabla V(x)$. This leads to a reformulation of the equation as:
$\nabla V(x)=-\lambda\dot{x}+\eta,$ (B.2)
or equivalently,
$\lambda\dot{x}=-\nabla V(x)+\eta.$ (B.3)
This relationship allows us to interpret the motion of the particle in terms of potential energy changes, where $-\nabla V(x)$ represents the deterministic force derived from the potential energy landscape and $\eta$ accounts for stochastic thermal fluctuations. To capture these dynamics in a more mathematically rigorous framework suitable for simulation and analysis, we can transform this physical concept into a stochastic differential equation (SDE):
$dx=-\nabla V(x)dt+\sqrt{2}\sigma dW,$ (B.4)
where the term $-\nabla V(x)$ is the drift term, which guides the particle towards lower potential energy states, thus simulating deterministic motion under the influence of forces. The term $\sqrt{2}\sigma dW$ represents the stochastic component of the motion, with $\sigma$ being the volatility (akin to the temperature in physical systems) and $dW$ denoting the derivative of a standard Wiener process, which models the random thermal fluctuations.
In the context of probability distributions, the stochastic process described in Equation (B.4) corresponds to a Fokker-Planck equation (Fokker, 1914; Planck, 1917; Kadanoff, 2000) for the probability density function $p(x,t)$ of finding the system in state $x$ at time $t$:
$\frac{\partial}{\partial t}p(x,t)=\nabla\cdot\left[\nabla V(x)p(x,t)+\sigma^{2}\nabla p(x,t)\right].$ (B.5)
At steady state, the time derivative of the probability density function becomes zero, implying:
$0=\nabla\cdot\left[\nabla V(x)p(x,t)+\sigma^{2}\nabla p(x,t)\right],$ (B.6)
which simplifies to:
$\nabla V(x)p(x)=-\sigma^{2}\nabla p(x,t).$ (B.7)
Dividing both sides by $p(x)$ and rearranging gives:
$\nabla V(x)\propto-\frac{\nabla p(x,t)}{p(x)}=-\nabla\log p(x),$ (B.8)
indicating that, at equilibrium, the gradient of the potential energy is proportional to the gradient of the log of the probability distribution. This equilibrium condition underlies the principle that in a score-based generative model, the score function $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ can be interpreted as akin to a force derived from a potential energy landscape, guiding the generation process towards high-probability regions of the data distribution.
Therefore, Equation (B.4) can be re-written as:
$dx=\nabla\log p(x)dt+\sqrt{2}\sigma dW,$ (B.9)
To implement this in a discrete setting for numerical simulation or data generation, the SDE can be approximated as follows:
$\mathbf{x}_{i+1}\leftarrow\mathbf{x}_{i}+\nabla_{x}\log p(\mathbf{x})\epsilon+\sqrt{2}\sigma\Delta W,$ (B.10)
where $\epsilon$ is a small timestep, and $\Delta W\sim\mathcal{N}(0,\epsilon)$ represents a discrete approximation of the Wiener process increment. Thus, it can be further simplified as:
$\mathbf{x}_{i+1}\leftarrow\mathbf{x}_{i}+\nabla_{x}\log p(\mathbf{x})\cdot\epsilon+\sqrt{2\epsilon}\cdot\mathbf{z}_{i},$ (B.11)
where $\mathbf{z}_{i}\sim\mathcal{N}(0,I)$.